### v1

# BhashaVerse : Translation Ecosystem for Indian Subcontinent Languages

12024
[…]
BhashaVerse : Translation Ecosystem for Indian Subcontinent Languages
issue: 1
1 Introduction
Generative Large Language Models (LLMs) have brought transformative advancements to natural language processing (NLP), excelling in a wide range of applications Xuanfan and Piji (2023); Xi et al. (2023). These models demonstrate remarkable capabilities in open-domain question answering by generating accurate and coherent responses and performing instruction-based tasks such as code completion, error detection, and correction Vaithilingam, Zhang, and Glassman (2022). Additionally, LLMs are highly effective in tasks like essay writing, grammar correction Wu et al. (2023), and text summarization, producing outputs of exceptional quality Chang et al. (2023). However, much of this progress has been concentrated on English, leaving many other languages underexplored or underdeveloped Lai et al. (2023); Zhu et al. (2023). For machine translation (MT) and related subtasks, encoder-decoder-based models remain state-of-the-art, while decoder-only models require further advancement to effectively handle low-resource languages, particularly those spoken in India.
India’s linguistic diversity is immense, comprising 22 scheduled languages from linguistic families such as Indo-Aryan, Dravidian, Tibeto-Burman, and Austroasiatic. There are over 559 classified mother tongues spoken across the country, reflecting the cultural and linguistic richness of the nation [^1] https://en.wikipedia.org/wiki/Linguistic˙Survey˙of˙India. This diversity, however, poses significant challenges in communication, education, business, healthcare, tourism, and governance. Advances in encoder-decoder and LLM-based models hold the potential to address these linguistic challenges by enabling cross-lingual communication and fostering India’s multilingual ecosystem.
Translation plays a vital role in addressing linguistic diversity and enabling effective cross-lingual communication in multilingual societies like India. However, the distinct characteristics of Indian languages present unique challenges. Many Indian languages exhibit complex morphological structures, often involving agglutinative or inflectional morphology, which makes them syntactically diverse and distinct from each other and English. The use of diverse scripts, such as Devanagari, Tamil, Bengali, and Gurmukhi, adds further complexity to text processing tasks. Additionally, code-mixing, a common phenomenon in India where speakers frequently blend multiple languages in communication, complicates translation workflows. The scarcity of high-quality linguistic resources, annotated corpora, and evaluation benchmarks for many Indian languages exacerbates these challenges, hindering the development of robust MT systems.
Addressing these challenges requires a systematic approach. Developing translation systems for Indian languages is crucial to facilitate communication and preserve linguistic diversity. With the translation system, robust evaluation frameworks are needed to assess their quality, with or without reference translations. Identifying and categorizing translation errors is essential for system improvement, enabling targeted refinements based on predefined error groups. Automatic post-editing systems can then help correct translation errors efficiently, either independently or in collaboration with human translators. Incorporating these advancements across general and domain-specific applications will greatly benefit the language research community and end-users.
This paper focuses on developing translation models and related applications for 36 Indian languages, including Assamese, Awadhi, Bengali, Bhojpuri, Braj, Bodo, Dogri, English, Konkani, Gondi, Gujarati, Hindi, Hinglish, Ho, Kannada, Kangri, Kashmiri (Arabic and Devanagari), Khasi, Mizo, Magahi, Maithili, Malayalam, Marathi, Manipuri (Bengali and Meitei), Nepali, Oriya, Punjabi, Sanskrit, Santali, Sinhala, Sindhi (Arabic and Devanagari), Tamil, Tulu, Telugu, and Urdu. Achieving this requires parallel and other types of corpora for all 36 × 36 language pairs, addressing challenges like script variations, phonetic differences, and syntactic diversity. For instance, languages like Kashmiri and Sindhi, which use multiple scripts, demand script normalization for alignment, while low-resource languages such as Khasi and Santali require synthetic data augmentation to ensure sufficient coverage and quality.
To address these challenges, this work proposes strategies for corpus creation by leveraging existing resources, developing parallel datasets, generating domain-specific corpora, and utilizing synthetic data techniques. Additionally, it evaluates machine translation across various dimensions, including standard and discourse-level translation, domain-specific translation, reference-based and reference-free evaluation, error analysis, and automatic post-editing. By integrating these elements, the study establishes a comprehensive framework to improve machine translation quality and enable better cross-lingual communication in India’s linguistically diverse ecosystem.
Finally, this work addresses several translation-related applications that are central to enhancing multilingual communication.
  1. 1.
    Translation Across 36 × 36 Indian Language Pairs: We develop machine translation models capable of translation for 36 Indian languages to 36 Indian languages. This involves creating n-way parallel corpora, aligning language pairs, and optimizing translation quality to meet linguistic and cultural expectations.
  2. 2.
    Discourse Translation: Translation at the discourse level addresses coherence, context retention, and inter-sentence dependencies. Our work evaluates how well current systems handle these aspects and proposes improvements for translating connected text effectively. Discourse translation possible for 3636 language pairs with keeping 1536 tokens as a context with encoder-decoder model while 4096 context for decoder only model.
  3. 3.
    Domain-Specific Translation: Different domains, such as healthcare, legal, and education, require specialized vocabulary and context-aware translations. We design systems capable of handling domain-specific language nuances to improve usability in practical scenarios. We include domains
  4. 4.
    Machine Translation Evaluation: We introduce both reference-based and reference-free evaluation methods to assess translation quality. While reference-based methods compare MT output against human-translated gold standards, reference-free approaches evaluate fluency, adequacy, and fidelity without explicit references.
  5. 5.
    Translation Error Identification and Categorization: To identify and address translation errors, we propose a taxonomy for error analysis that includes lexical, grammatical, semantic, and discourse-level errors. This categorization aids in pinpointing system weaknesses and guiding future improvements.
  6. 6.
    Automatic Post-Editing (APE): Post-editing involves correcting errors in machine-generated translations to improve fluency and accuracy. We develop automatic post-editing systems tailored for Indian languages, leveraging error patterns to refine translations.
By addressing these interconnected challenges, our work aims to advance the state of machine translation for Indian languages, bridging linguistic gaps and fostering multilingual communication in diverse settings. The remainder of this paper is organized as follows: Section 2 presents related work in machine translation and evaluation for low-resource languages, with a focus on Indian languages. Section 3 details the data collection and preprocessing methods used for building parallel corpora and specialized datasets. Section 4 describes the translation models and architectures employed for sentence, discourse, and domain translation tasks. Section 5 elaborates on the evaluation framework, including metrics and human assessment strategies. Section 6 introduces our taxonomy for error identification and discusses the implementation of automatic post-editing systems. Section 7 presents experimental results, analysis, and comparisons with existing systems. Section 8 concludes the paper with a discussion of findings and future research directions.
2 Corpora : Existing Machine Translation Corpora
Corpora are essential for developing language processing systems, especially for low-resource, morphologically complex languages like those in the Indian subcontinent. They underpin the training, evaluation, and tuning of machine translation (MT) models, ensuring diverse and accurate outputs. This section explores existing and newly developed corpora for MT, focusing on preparation, cleaning, alignment, and generation. Indian languages pose unique corpus creation challenges due to script, grammar, and cultural diversity. We use a mix of curated parallel corpora, domain-specific datasets, and synthetic data to address these, facilitating robust translation models for 36 Indian languages and their combinations.
Machine translation (MT) research for Indian languages is heavily reliant on robust parallel corpora that address the unique linguistic, resource, and domain-specific challenges inherent in India’s multilingual landscape. Over the years, a variety of datasets have been developed to support translation between Indian and other languages, as well as Indian-to-Indian language pairs, enabling significant progress in this field.
The BPCC Parallel Corpus (Gala et al., 2023) represents the largest publicly accessible resource, a collection of several existing and developed parallel corpora resources, consisting of 230 million sentence pairs across English and 22 Indian languages. It features the BPCC-Human Corpus, a refined subset containing 2.2 million English-Indic sentence pairs that have been manually validated, thereby providing dependable data for both training and evaluation purposes. In a similar context, the Samanantar Parallel Corpus (Ramesh et al., 2022) was, until 2021, the most extensive collection of parallel corpora for English and 11 Indian languages, offering 46 million English-Indic sentence pairs and 82 million Indian-to-Indian language pairs. This corpus serves as a crucial dataset for multilingual machine translation research.
Corpora such as the IIT Bombay English-Hindi Parallel Corpus (1.5 million segments) (Kunchukuttan, Mehta, and Bhattacharyya, 2017) and PMIndia (Haddow and Kirefu, 2020) provide domain-specific resources, particularly for English-Hindi and other English-Indic translations. PMIndia, derived from government texts like the Prime Minister’s Mann Ki Baat speeches (Philip et al., 2019), and the CVIT-IIITH PIB Multilingual Corpus (Philip et al., 2019), mined from Press Information Bureau [^2] https://pib.gov.in/ web pages, emphasize high-quality, formal translations. These datasets enrich MT research with domain-specific data critical for practical applications.
For Indian-to-Indian language translation, resources are more limited. The LTRC Parallel Corpus [^3] https://github.com/vmujadia/The-LTRC-Hindi-Telugu-Parallel-Corpus for Hindi-Telugu (Mujadia and Sharma, 2022), developed by IIIT Hyderabad, offers 500,000 sentence-aligned translations, providing robust data for these structurally distinct languages. Another work focusing on a low-resource language pair, Hindi-Kangri, presented a corpus (Chauhan, Saxena, and Daniel, 2021) with 27,000 parallel sentences. The IndoWordNet Parallel Corpus (Bhattacharyya, 2020), with 6.3 million segments across 18 languages, is another key resource, leveraging glosses and examples to facilitate Indian-to-Indian language MT development.
Low-resource languages are addressed through specialized corpora such as the Ema-lon Manipuri Corpus (124,975 Manipuri-English aligned sentences) (Huidrom, Lepage, and Khomdram, 2021) and NLLB-Seed [^4] https://github.com/facebookresearch/flores/blob/main/nllb˙seed/README.md, https://github.com/openlanguagedata/seed, which focuses on low resourced languages like Kashmiri, Maithili, and Bhojpuri among others. These efforts aim to provide critical data for languages often underrepresented in MT research. Code-mixed languages, a common phenomenon in India, are supported by datasets like PHINC (Srivastava and Singh, 2020) (Hinglish-English, 13,738 sentences) and the IIIT-H en-hi-codemixed-corpus (Dhar, Kumar, and Shrivastava, 2018) (6,096 English-Hindi code-mixed sentences), which are essential for handling multilingual and code-switched input.
Domain-specific corpora further enhance the scope of Indian languages machine translation, Itihasa Parallel Corpus (Aralikatte et al., 2021) offers 93,000 parallel sentences for English-Sanskrit translations, catering to ancient language research, while the TED Parallel Corpus [^5] https://github.com/ajinkyakulkarni14/TED-Multilingual-Parallel-Corpus and JW300 Corpus (Agic and Vulic, 2019) provide resources for public speeches and religious texts. Regional languages are supported by corpora like CGNetSwara [^6] http://cgnetswara.org/ Hindi-Gondi (19,000 sentence pairs) and MTEnglish2Odia [^7] https://github.com/soumendrak/MTEnglish2Odia (42,000 pairs), filling critical gaps for underrepresented groups.
Large-scale mining initiatives like CCAligned (100 million document pairs across 137 languages)(El-Kishky et al., 2019) and CoPara (Nikhil, Choudhary, and Mamidi, 2023)(passage-level alignments for Dravidian languages) significantly expand multilingual datasets. While noisy, resources like WikiMatrix(Schwenk et al., 2019a) and CCMatrix (Schwenk et al., 2019b) mined from Wikipedia and CommonCrawl offer valuable data for pre-processing pipelines. Other notable resources include the Nepali National Corpus (Yadava et al., 2008), the Kathmandu University English-Nepali Parallel Corpus (Duwal and Bal, 2019), and the Uka Tarsadia University Corpus for English-Gujarati translations (Shah and Bakrola, 2019), each addressing specific regional and linguistic needs.
The Low-Resource MT Shared Task  [^8] https://sites.google.com/view/loresmt/  [^9] https://www.statmt.org/wmt21/similar.html  [^10] https://github.com/loresmt Corpus, used in platforms like the Workshop on Asian Translation (WAT)  [^11] https://lotus.kuee.kyoto-u.ac.jp/WAT/WAT2024/index.html, supports MT systems for Hindi and other low-resource languages. These datasets combine Wikipedia, government documents, and religious texts with crowd-sourced translations, focusing on language pairs like Hindi-Marathi, Hindi-Telugu, Tamil-Telugu, and Hindi-Bhojpuri. Synthetic techniques like back-translation (Gala et al., 2023) and multilingual pretraining are increasingly used to augment these corpora, improving model performance for low-resource languages. These efforts drive innovation in transfer learning, multilingual modeling, and domain-specific translation, addressing critical gaps in India’s MT landscape.
Collectively, these corpora form a robust foundation for machine translation research and development work for Indian languages, advancing cross-lingual communication and supporting the preservation of India’s linguistic diversity.
3 Developed Corpora : Machine Translation
In this section, we detail the methodologies that we used to build corpora for training machine translation (MT) systems, focusing on Indian subcontinent languages. We explore alignment techniques, domain-specific corpora development, synthetic data generation methods, and corpus cleaning methodologies to build more than 1B parallel corpora for machine translation involving 36 Indian subcontinent languages.
3.1 Automatic Alignment and Human Validation
We collected data from various multilingual websites and books containing content in English and multiple Indian languages. Using a neural alignment tool [^12] https://github.com/vmujadia/sentencealigner based on the COMET-QE model (Rei et al., 2022b), we calculated sentence-level similarity scores to align multilingual webpage content. This allowed us to identify English sentences and their corresponding translations in other languages, facilitating the creation of parallel corpora. For each language pair, specific alignment thresholds131313determined based on the average COMET-QE score for the respective language pair were applied, enabling the extraction of aligned data for pairs such as English-Assamese, English-Bangla, English-Gujarati, English-Hindi, English-Kannada, English-Malayalam, English-Marathi, English-Odia, English-Punjabi, English-Tamil, and English-Telugu.
The process presented challenges due to the diverse sentence structures, significant word order variations, and rich morphological features inherent to Indian languages. To address these, we conducted human validation on 10% of the sampled data to assess and adjust the alignment thresholds, ensuring the creation of high-quality parallel corpora suitable for training machine translation systems.
3.2 Human Post-edited Education and Medical Domain Parallel Corpora
Domain-specific corpora are essential for developing machine translation (MT) systems that perform exceptionally well in specialized areas such as medicine and technical content. These corpora capture critical domain-specific terminology, language expressions, and contextual nuances, disfluencies managed in translation if the content is spoken, which are often missing from general-purpose datasets.
To build parallel corpora, we enlisted multiple freelance translation experts to translate text from English into Indian languages, including Assamese, Gujarati, Hindi, Kannada, Malayalam, Marathi, Odia, Tamil, and Telugu. The primary focus was on educational domains, utilizing text collected from technical lectures, including those sourced from the Swayam [^14] https://swayam.gov.in/ and NPTEL [^15] https://nptel.ac.in/ platforms. For medical data, the corpus included content from various medical research protocols for research studies, such as consent forms, information sheets, and medical awareness materials. These were obtained from a well-known private Christian minority community-run medical college and hospital. We first transcribed these video contents verbatim with the help of respective language experts, ensuring accurate transcription. Subsequently, post-editing was performed on machine-translated content generated using multiple translation engines, such as SSMT [^16] http://ssmt.iiit.ac.in/translate and Google Translate [^17] https://translate.google.co.in/. Given the technical nature of the task, we provided clear guidelines to the translators to ensure consistent quality and adherence to domain standards. The guidelines included:
  • Thorough Understanding: Translators need to carefully read the entire source text before beginning the translation process to grasp its meaning and context fully.
  • Faithfulness to the Source Translations need to accurately capture the meaning of each sentence and remain as clear and understandable as the original text as a whole.
  • Rigorous Review It is advised to review own work multiple times, both silently and aloud, to ensure clarity, coherence, and appropriateness of word choice.
  • Consistency in Terminology Technical terms and expressions are have to be used consistently throughout the translation. Translators need to rely on established term translations present in resources like NCERT textbooks [^18] https://ncert.nic.in/textbook.php. If a term lacked a clear translation, it is to be transliterated.
  • Contextual Alignment The overall translation has to convey the same natural message and context as the original text.
To ensure quality, an in-house team of experienced translators (4 to 5 experts for each language pair) rigorously validated each sentence translated by freelancers following the same guidelines. Discrepancies or deviations from the guidelines are corrected by reviewers as part of the process. All these activities are conducted using a custom-built collaborative language platform, PostEditMe [^19] https://posteditme.in/, which effectively managed workflows for translation, post-editing, validation, and task assignments throughout this effort. The developed human post-edited parallel corpora contain healthcare and educational domains spanning over a wide range of sub-domains, including Psychology, Natural Sciences, Teaching Methods, Mathematics, Political Science, History, Communication Skills, Law, Tourism, Computer Science, Marketing, Management, Textile, Economics, and Health, ensuring comprehensive coverage of diverse academic disciplines. Notably, the post-edited parallel corpora were developed with English as the source language, and the same English text was post-edited into other Indian languages, enabling the creation of n-way parallel corpora among Indian languages.
3.3 Quality Synthetic Parallel Corpora
Synthetic corpora can address the paucity of parallel data for Indian languages. To bridge this gap, we utilized various techniques, including pivot-based translation, back-translation, and forward translation at the sentence and paragraph level. These methods significantly enhance the scope of existing corpora, expanding their coverage and offering vital support for low-resource language pairs.
3.3.1 Pivoted Parallel Corpora
Most of the parallel corpora developed for Indian languages focus on translations involving English, with relatively few efforts dedicated to creating corpora for Indian-to-Indian language translations. However, such corpora are crucial for addressing the linguistic diversity and translation needs of the Indian subcontinent. Pivot-based translation offers an effective solution to the lack of direct parallel corpora for underrepresented language pairs.
In this work, we enrich Indian-to-Indian language corpora by generating high-quality synthetic data using pivot translation. Our approach uses well-represented languages such as English and Hindi as intermediary "pivot" languages to bridge the gap between other low-resource Indian languages. We leveraged existing corpora for these languages and employed SSMT Translator [^20] http://ssmt.iiit.ac.in/translatev3 to translate from Indian languages to English, and then from English to other Indian languages under consideration. Similarly, translations are generated between Indian languages and Hindi, using Hindi as a pivot. This process involves creating automatic translations first between Indian languages and the pivot (English or Hindi), followed by translations from the pivot to the target Indian languages.
To address the varying quality of generated parallel corpora, we implemented filtering mechanisms to ensure the reliability of the training data. Advanced methods, such as COMET-QE (Rei et al., 2022a), were employed to evaluate the quality of sentence pairs by assigning quality estimation scores. Sentence pairs falling below a set threshold (determined as the average QE score across all pairs for each language pair) were filtered out, maintaining a high standard for the training datasets. To further validate this approach, randomly selected sentences from the filtered corpora were reviewed by in-house language experts. Their evaluations confirmed that the majority of the automatically scored and filtered data were valid parallel corpora, demonstrating its effectiveness in supporting machine translation models and enhancing their performance and generalizability. By leveraging existing translations involving the pivot languages, this method enables the generation of new parallel corpora, thereby extending the reach of machine translation models to low-resource language pairs. This approach facilitates cross-lingual applications and fosters inclusivity in multilingual natural language processing.
3.3.2 Iterative Backward Translation
We employed back-translation, a widely used data augmentation technique in machine translation, to enhance the diversity and quality of training data. Monolingual corpora for the target languages were sourced from WMT News Crawl [^21] https://data.statmt.org/news-crawl/ and Wikipedia [^22] https://huggingface.co/datasets/wikimedia/wikipedia. Using SSMT Translator [^23] http://ssmt.iiit.ac.in/translatev3 and IndicTrans2242424https://github.com/AI4Bharat/IndicTrans2, we translated this monolingual data bidirectionally: from Indian languages to English and from English to Indian languages. The translated data introduced variations in sentence structure, vocabulary, and phrasing, enriching the training set’s diversity. For each monolingual corpus, both translation outputs were evaluated using COMET-QE scores, and the version with the highest score was retained. This filtered data was then used to train a multilingual machine translation model involving the selected language pairs. The newly trained model was subsequently used to re-translate the same monolingual corpora, iteratively improving the translations. This process was repeated for five iterations, with the best-scored translation pairs from all iterations being selected for the final training corpus.
Paragraph Level Backward Translation
For the Wikipedia dataset [^25] https://huggingface.co/datasets/wikimedia/wikipedia and crawled news articles from the internet, we adopted paragraph-level backward translation to preserve coherence and contextual consistency across entire paragraphs. Unlike sentence-level back-translation, this approach ensures that the translated content retains discourse-level attributes, such as topic continuity, pronoun resolution, and lexical cohesion. By focusing on paragraph-level fidelity, this method enhances the quality of machine translation systems, particularly for tasks requiring deeper contextual understanding, such as document translation or conversational AI. The same filtering mechanism was applied here as well. Each sentence within a paragraph was evaluated for quality using COMET-QE scores, and the best-scored translations were selected. These high-quality sentence translations were then reassembled into paragraphs to maintain the original discourse structure. This strategy ensured the development of a robust multilingual machine translation model capable of handling complex and contextually rich translations.
3.4 Parallel Corpora Cleaning
Cleaning parallel corpora is crucial for ensuring the accuracy and reliability of training datasets. Noise in parallel corpora, such as duplicates, incomplete translations, misalignments, script variations, and unwanted symbols, can significantly degrade the performance of machine translation (MT) models. To address these issues, we utilized automated tools to detect and resolve errors or remove problematic sentence pairs. This process involved normalizing text and tackling script-specific challenges in Indian languages, such as script diversity and spelling inconsistencies. Additional preprocessing steps included removing mistranslations, addressing code-mixing, and correcting tokenization errors. We employed multiple methodologies to achieve effective corpus cleaning:
Length-Based Filtering
This method filters parallel corpora by comparing differences in word and character counts to ensure alignment quality. For example, we calculate the average sentence length for each language and maintain an acceptable range of differences for a given language pair in terms of words and characters. This range, set as the threshold, allows for a variation of ±10 words (or corresponding characters). Every source and target text pair in the generated corpora is evaluated against this threshold. Sentence pairs that do not meet this criterion are discarded to maintain the quality of the dataset.
Language and Script Identification
We utilized a FastText-based language detection tool to identify the language of each word within a sentence. If the majority of word-level language tags do not align with the sentence-level language tag, the sentence pair is filtered out to ensure linguistic consistency.
HTML/XML Tag Validation
Sentence pairs with mismatched HTML or XML tags between the source and target sides are removed to avoid structural inconsistencies in the data.
COMET-QE Scoring
The quality of source-target and target-source sentence pairs was evaluated using COMET-22 scores. Sentence pairs with scores significantly below the average threshold (average minus 10) were excluded, ensuring that only high-quality data was retained for training.
We applied these cleaning methods to existing, newly developed, and synthetically generated parallel corpora for Indian languages to achieve greater alignment accuracy, noise reduction, and linguistic consistency, resulting in more effective MT systems.
4 Automatic Post-Editing (APE) Corpora
We propose methods to enhance corpora for automatic post-editing between English and Indian languages. These include using human post-edited corpora and generating synthetic corpora from existing parallel data, providing a foundation for general and domain-specific post-editing tasks.
4.1 Human APE Corpora
As discussed in Section 3.2, post-editing involves human translators refining machine-generated translations to produce accurate and natural-sounding final translations. This process was conducted using the PostEditMe [^26] https://posteditme.in/ workbench, a platform designed to facilitate post-editing tasks. The workbench allowed translators to edit machine-generated translations while maintaining a record of both the original machine-generated version and the human-corrected version. Additionally, the workbench captured all intermediate stages of post-editing, recording multiple iterations for each sentence pair, along with the final edited version. These intermediate stages and the final post-edited text form a comprehensive resource for studying the post-editing process and its impact on translation quality. The resulting dataset, named the Bhashik APE corpus, includes the initial machine-generated (also intermediate stages text) and human-post-edited translations for source texts in English and 11 Indian languages. This corpus supports the development of automatic post-editing systems and enhances domain-specific machine translation by addressing the linguistic nuances of technical and educational contexts more effectively.
4.2 Synthetically Developed Corpora
Automatic post-editing (APE) corpora typically consist of three key components: the source text, an imperfect or erroneous translation, and the error-free translation. To synthetically generate such APE corpora, we utilized both existing parallel corpora and our newly developed parallel corpora described in this work. The target sentences or passages are subjected to various perturbation techniques as detailed in Appendix Section 7, simulating errors commonly found in machine translations. For our opted language pairs, we processed each sentence pair to create multiple denoised versions of the target text. These perturbations introduced varying degrees of errors, ranging from 2% to 15% of the total words, including grammar mistakes, mistranslations, and word order issues. These artificially generated error patterns simulate typical MT errors, making the perturbed version function as the erroneous translation, while the original target text serves as the error-free translation. Additionally, we applied noisy back-translation techniques to create diverse post-editing pairs, ensuring a wide range of error types for robust APE training data. This approach provides a rich and diverse dataset for training automatic post-editing models. We call the resulting dataset Bhashik APE (Synth).
5 Corpora for Machine Translation Evaluation
In this work, we developed machine translation assessment corpora for English and Indian languages through multiple methods to evaluate translation quality. We focus on translation evaluation which is on a continuous scale from 1 to 100, where 1 indicates poor translation and 100 represents the best possible translation.
We use two approaches: reference-based evaluation, comparing translations to a reference, and reference-free evaluation, assessing translations without references. Corpora development involves human expert scoring and synthetic data generation using perturbation techniques (Appendix Section 7) to simulate common translation errors. Human assessment corpora are created by experts scoring translations on a continuous scale, while synthetic data is generated with COMET models scoring translations across scenarios derived from parallel corpora. This section outlines the strategies for creating synthetic corpora to comprehensively evaluate machine translation systems.
5.1 Human Annotated Corpora
Human-annotated corpora are created by language experts who evaluate translations based on fluency, adequacy, and contextual accuracy, providing a gold standard for assessing MT systems. Experts fluent in both the source and target languages assess the accuracy of the translation in conveying the original meaning, focusing on Translation Accuracy, which determines whether the translation preserves the source text’s meaning while maintaining fluency and grammatical correctness. Scores are assigned on a scale of 1 to 100.
To develop these corpora, we utilized benchmark datasets and multiple translation engines, including Google Translate, SSMTv3, and Indictransv2, to generate translations for input texts. Experts evaluated both the source text and the translated text, providing ratings. The reference translations in the benchmark corpora were used for reference-enabled machine translation evaluation, while the annotations can also be applied in reference-free evaluation scenarios. The dataset spans 18 language pairs and 36 translation directions, resulting in over 72,000 rated translations. This effort involved over 70 experienced language experts, with each source and target translation pair receiving three independent ratings.
5.2 Synthetically Developed Corpora
To achieve comprehensive coverage of the translation evaluation spectrum on a linear scale, we recognized that relying solely on existing or limited human scored direct assessment data is insufficient. Therefore, we decided to generate additional data synthetically using existing parallel and benchmark corpora. Various techniques are employed to create degraded, perturbed versions of translations, and we used an existing COMET model to assign degraded scores using the following methodology:
  1. 1.
    Generating Perturbed Translations: We began by sourcing parallel and benchmark corpora from above mentioned repositories. For each sentence pair, we applied perturbation techniques outlined in Appendix Section 7 to generate lower-quality translations with varying degrees of errors. The error levels ranged from 5% to 50%.
  2. 2.
    Structuring Sentence Tuples: For each perturbed version, we created a sentence tuple containing the following elements:
    • Source text (S)
    • Original translation text (T) from the parallel corpora
    • Perturbed translation text (T̃) with Error percentage (Ep)
  3. 3.
    Calculating Synthetic Scores: Using the existing COMET Quality Estimation (QE) model, we first evaluated the source text (S) and the original translation (T) to obtain a COMET score (Cs) and 100 minus the Translation Error Rate (TER) score (Tp) using original translation (T) and Perturbed translation text (T̃). We then averaged the error percentage (Ep) and the TER-derived score (Tp) to calculate a degradation factor. This factor was subtracted from the COMET score (Cs) to derive a new synthetic score for the source text (S) and the perturbed translation (T̃).
By applying this process, we generated an extensive synthetic dataset to supplement the limited human-annotated data, enabling robust evaluation across a broader range of translation quality scenarios. The resulting dataset is named the Bhashik PseudoMTDA corpus for English and 11 Indian languages.
6 Corpora for Machine Translation Error Span Identification and Error Category Marking
Error identification and categorization help analyze weaknesses in MT systems, enabling targeted improvements. This section outlines our development of corpora to support these tasks.
6.1 Human MT Error Span Identification Corpora
As discussed in Section 3.2, post-editing involves human translators refining machine-generated translations to produce accurate, natural-sounding final translations. This process was carried out using the PostEditMe [^27] https://posteditme.in/ workbench, a platform specifically designed to support post-editing tasks. The workbench captures every edit made by translation experts and records this information at the word level over the machine translation output. The edited spans in the translated text, where changes were made, are marked as MT error spans for the corresponding translation. These post-editing efforts provide a rich resource for identifying error spans and evaluating translation quality. The resulting dataset, named the Bhashik PseudoMTError corpus, includes the source text, the initial machine-generated translation, and the annotated error spans for English and 11 Indian languages.
6.2 Synthetically Developed MT Error Span and Category Marking Corpora
As discussed earlier, corpora for machine translation (MT) error span and category marking typically include three key components: the source text, a translation annotated with error spans, and corresponding error categories. To synthetically create such corpora, we leveraged both existing parallel corpora and newly developed parallel corpora introduced in this work. Target sentences or passages were subjected to various perturbation techniques, detailed in Appendix Section 7, to simulate common machine translation errors across different categories. These categories were aligned with broader MQM categories.
For the selected language pairs, we processed each sentence pair to generate multiple denoised versions of the target text. Perturbations introduced errors ranging from 2% to 15% of the total words, including grammar mistakes, mistranslations, and word order issues. These artificially generated error patterns mimic typical MT errors, making the perturbed versions serve as erroneous translations. Error spans were annotated based on the introduced errors, along with the corresponding categories used to generate the perturbations.
This approach provides a rich and diverse dataset for training automatic systems to mark MT error spans and categories. We refer to the resulting dataset as the Bhashik PseudoMTError Corpus.
6.3 Developed Parallel Corpora
6.3.1 Bhashik Education and Medical: Human Post-edited Corpora
Table 1: The post-edited 2M domain-parallel corpora spans English and 10 Indian languages, covering 15 educational domains.
Domain Language Pair Passages/Sentences Domain Language Pair Passages/Sentences
Natural science eng_Latn-hin_Deva 140999 Teaching methods eng_Latn-tam_Taml 17
Natural science eng_Latn-mar_Deva 191936 Teaching methods eng_Latn-mal_Mlym 527
Natural science eng_Latn-kan_Knda 185491 Teaching methods eng_Latn-urd_Arab 3909
Natural science eng_Latn-tel_Telu 223146 Management eng_Latn-mar_Deva 14521
Natural science eng_Latn-mal_Mlym 84 Management eng_Latn-kan_Knda 11001
Natural science eng_Latn-tam_Taml 38 Management eng_Latn-tel_Telu 11041
Computer science eng_Latn-mar_Deva 80170 Health eng_Latn-hin_Deva 715
Computer science eng_Latn-hin_Deva 71507 Psychology eng_Latn-hin_Deva 12651
Computer science eng_Latn-kan_Knda 75354 Psychology eng_Latn-tel_Telu 12771
Computer science eng_Latn-tel_Telu 75388 Psychology eng_Latn-mar_Deva 12822
Computer science eng_Latn-urd_Arab 332 Psychology eng_Latn-kan_Knda 12570
Computer science eng_Latn-tam_Taml 11902 Marketing eng_Latn-hin_Deva 4903
Law eng_Latn-tel_Telu 80165 Marketing eng_Latn-tel_Telu 4971
Law eng_Latn-mar_Deva 81581 Marketing eng_Latn-mar_Deva 4858
Law eng_Latn-tam_Taml 90 Marketing eng_Latn-kan_Knda 5099
Law eng_Latn-kan_Knda 78245 Political science eng_Latn-tel_Telu 11978
Law eng_Latn-hin_Deva 735 Political science eng_Latn-mar_Deva 11935
Mathematics eng_Latn-mar_Deva 62881 Political science eng_Latn-kan_Knda 12001
Mathematics eng_Latn-tel_Telu 59373 Political science eng_Latn-hin_Deva 7876
Mathematics eng_Latn-kan_Knda 40625 Textile eng_Latn-tel_Telu 7609
Mathematics eng_Latn-hin_Deva 26259 Textile eng_Latn-mar_Deva 7637
Mathematics eng_Latn-tam_Taml 286 Textile eng_Latn-kan_Knda 7630
Mathematics eng_Latn-mal_Mlym 163 History eng_Latn-tel_Telu 4902
health eng_Latn-tam_Taml 12728 History eng_Latn-mar_Deva 4919
health eng_Latn-guj_Gujr 299 History eng_Latn-kan_Knda 4941
health eng_Latn-tel_Telu 34855 Economics eng_Latn-tel_Telu 5087
health eng_Latn-hin_Deva 11802 Economics eng_Latn-mar_Deva 5093
health eng_Latn-mar_Deva 1904 Economics eng_Latn-kan_Knda 5061
health eng_Latn-kan_Knda 135 Tourism eng_Latn-tel_Telu 12798
health eng_Latn-mal_Mlym 202 Tourism eng_Latn-mar_Deva 12795
health eng_Latn-ben_Beng 370 Tourism eng_Latn-kan_Knda 12813
Teaching methods eng_Latn-kan_Knda 32670 Communication skill eng_Latn-tel_Telu 8426
Teaching methods eng_Latn-tel_Telu 31828 Communication skill eng_Latn-mar_Deva 8537
Teaching methods eng_Latn-hin_Deva 19065 Communication skill eng_Latn-kan_Knda 8627
Teaching methods eng_Latn-mar_Deva 32327
6.3.2 Bhashik Generic : Quality Synthetic Corpora
Table 2: The generated 1B parallel corpora cover 325 language pairs, encompassing 36 languages from the Indian subcontinent and utilizing 15 unique scripts.
Language Pair
Developed
Parallel Data
Language Pair
Developed
Parallel Data
Language Pair
Developed
Parallel Data
Language Pair
Developed
Parallel Data
Language Pair
Developed
Parallel Data
Language Pair
Developed
Parallel Data
Language Pair
Developed
Parallel Data
asm_Beng_ben_Beng 2111717 brx_Deva_eng_Latn 21161986 eng_Latn_mai_Deva 25730422 guj_Gujr_snd_Deva 7948035 kas_Arab_snd_Arab 872206 mar_Deva_ory_Orya 47036189 pan_Guru_snd_Deva 3246797
asm_Beng_brx_Deva 765615 brx_Deva_gom_Deva 799906 eng_Latn_mal_Mlym 103881922 guj_Gujr_tam_Taml 69922990 kas_Arab_snd_Deva 1231344 mar_Deva_pan_Guru 4421324 pan_Guru_tam_Taml 4207666
asm_Beng_doi_Deva 2317597 brx_Deva_guj_Gujr 3423331 eng_Latn_mar_Deva 98133618 guj_Gujr_tel_Telu 635961470 kas_Arab_tam_Taml 2359634 mar_Deva_san_Deva 1940249 pan_Guru_tel_Telu 80308819
asm_Beng_eng_Latn 30594977 brx_Deva_hin_Deva 7883998 eng_Latn_mni_Beng 13473789 guj_Gujr_urd_Arab 171350512 kas_Arab_tel_Telu 10578313 mar_Deva_sat_Olck 539452 pan_Guru_urd_Arab 7873795
asm_Beng_gom_Deva 828224 brx_Deva_kan_Knda 2247979 eng_Latn_mni_Mtei 24917741 hin_Deva_kan_Knda 27641372 kas_Arab_urd_Arab 3362396 mar_Deva_snd_Arab 2047746 san_Deva_sat_Olck 355047
asm_Beng_guj_Gujr 9163504 brx_Deva_kas_Arab 710060 eng_Latn_npi_Deva 79094284 hin_Deva_kas_Arab 10689877 kas_Deva_mai_Deva 1723270 mar_Deva_snd_Deva 2682147 san_Deva_snd_Arab 1135478
asm_Beng_hin_Deva 8361026 brx_Deva_kas_Deva 1429735 eng_Latn_ory_Orya 434864407 hin_Deva_kas_Deva 18466409 kas_Deva_mal_Mlym 3812953 mar_Deva_tam_Taml 3041800 san_Deva_snd_Deva 1453221
asm_Beng_kan_Knda 2374887 brx_Deva_mai_Deva 780823 eng_Latn_pan_Guru 106211275 hin_Deva_mai_Deva 8774341 kas_Deva_mar_Deva 3342699 mar_Deva_tel_Telu 79927241 san_Deva_tam_Taml 2085085
asm_Beng_kas_Arab 880736 brx_Deva_mal_Mlym 1986655 eng_Latn_san_Deva 43354388 hin_Deva_mal_Mlym 81376763 kas_Deva_mni_Beng 819979 mar_Deva_urd_Arab 40450921 san_Deva_tel_Telu 29210615
asm_Beng_kas_Deva 1577560 brx_Deva_mar_Deva 1462688 eng_Latn_sat_Olck 9625012 hin_Deva_mar_Deva 80444061 kas_Deva_mni_Mtei 1503829 mni_Beng_mni_Mtei 356775 san_Deva_urd_Arab 3636643
asm_Beng_mai_Deva 757165 brx_Deva_mni_Beng 378232 eng_Latn_snd_Arab 31294443 hin_Deva_mni_Beng 3728206 kas_Deva_npi_Deva 2457536 mni_Beng_npi_Deva 389988 sat_Olck_snd_Arab 348940
asm_Beng_mal_Mlym 1776146 brx_Deva_mni_Mtei 758003 eng_Latn_snd_Deva 31606531 hin_Deva_mni_Mtei 9416582 kas_Deva_ory_Orya 19983032 mni_Beng_ory_Orya 3585322 sat_Olck_snd_Deva 489121
asm_Beng_mar_Deva 1418576 brx_Deva_npi_Deva 954826 eng_Latn_tam_Taml 94940774 hin_Deva_npi_Deva 10786012 kas_Deva_pan_Guru 4201119 mni_Beng_pan_Guru 1119674 sat_Olck_tam_Taml 430418
asm_Beng_mni_Beng 312229 brx_Deva_ory_Orya 7615181 eng_Latn_tel_Telu 960415291 hin_Deva_ory_Orya 346421975 kas_Deva_san_Deva 1941751 mni_Beng_san_Deva 525579 sat_Olck_tel_Telu 2579655
asm_Beng_mni_Mtei 888899 brx_Deva_pan_Guru 2189039 eng_Latn_urd_Arab 246732366 hin_Deva_pan_Guru 63333450 kas_Deva_sat_Olck 676632 mni_Beng_sat_Olck 139996 sat_Olck_urd_Arab 976839
asm_Beng_npi_Deva 1009181 brx_Deva_san_Deva 751939 gom_Deva_guj_Gujr 9962240 hin_Deva_san_Deva 29108161 kas_Deva_snd_Arab 1837653 mni_Beng_snd_Arab 609824 snd_Arab_snd_Deva 1292311
asm_Beng_ory_Orya 12801566 brx_Deva_sat_Olck 240046 gom_Deva_hin_Deva 13998575 hin_Deva_sat_Olck 3377744 kas_Deva_snd_Deva 1827310 mni_Beng_snd_Deva 659939 snd_Arab_tam_Taml 1890070
asm_Beng_pan_Guru 2110174 brx_Deva_snd_Arab 900441 gom_Deva_kan_Knda 2432914 hin_Deva_snd_Arab 18584708 kas_Deva_tam_Taml 3329492 mni_Beng_tam_Taml 719910 snd_Arab_tel_Telu 19075895
asm_Beng_san_Deva 1034015 brx_Deva_snd_Deva 1060504 gom_Deva_kas_Arab 907722 hin_Deva_snd_Deva 14552624 kas_Deva_tel_Telu 19804719 mni_Beng_tel_Telu 2192230 snd_Arab_urd_Arab 3048720
asm_Beng_sat_Olck 439289 brx_Deva_tam_Taml 1321496 gom_Deva_kas_Deva 1739739 hin_Deva_tam_Taml 76095904 kas_Deva_urd_Arab 5387510 mni_Beng_urd_Arab 1254597 snd_Deva_tam_Taml 2661636
asm_Beng_snd_Arab 929278 brx_Deva_tel_Telu 6450144 gom_Deva_mai_Deva 760450 hin_Deva_tel_Telu 854298767 mai_Deva_mal_Mlym 2096537 mni_Mtei_npi_Deva 1443672 snd_Deva_tel_Telu 15779583
asm_Beng_snd_Deva 1315760 brx_Deva_urd_Arab 2250715 gom_Deva_mal_Mlym 2020062 hin_Deva_urd_Arab 211718893 mai_Deva_mar_Deva 1616046 mni_Mtei_ory_Orya 9403109 snd_Deva_urd_Arab 4222851
asm_Beng_tam_Taml 1645025 doi_Deva_eng_Latn 56050632 gom_Deva_mar_Deva 1393256 kan_Knda_kas_Arab 2720541 mai_Deva_mni_Beng 549939 mni_Mtei_pan_Guru 2435659 tam_Taml_tel_Telu 74713991
asm_Beng_tel_Telu 17575936 doi_Deva_gom_Deva 2230764 gom_Deva_mni_Beng 389990 kan_Knda_kas_Deva 4425169 mai_Deva_mni_Mtei 849017 mni_Mtei_san_Deva 1042586 tam_Taml_urd_Arab 22118281
asm_Beng_urd_Arab 2786265 doi_Deva_guj_Gujr 13475348 gom_Deva_mni_Mtei 1025298 kan_Knda_mai_Deva 2465368 mai_Deva_npi_Deva 1095355 mni_Mtei_sat_Olck 329603 tel_Telu_urd_Arab 242494781
ben_Beng_brx_Deva 1771330 doi_Deva_hin_Deva 25314841 gom_Deva_npi_Deva 1344302 kan_Knda_mal_Mlym 6077765 mai_Deva_ory_Orya 8857273 mni_Mtei_snd_Arab 1156456
ben_Beng_doi_Deva 5725017 doi_Deva_kan_Knda 6225786 gom_Deva_ory_Orya 11288994 kan_Knda_mar_Deva 4775780 mai_Deva_pan_Guru 2318762 mni_Mtei_snd_Deva 1302097
ben_Beng_eng_Latn 110423482 doi_Deva_kas_Arab 2260841 gom_Deva_pan_Guru 1984080 kan_Knda_mni_Beng 1409277 mai_Deva_san_Deva 1043353 mni_Mtei_tam_Taml 1748302
ben_Beng_gom_Deva 2152903 doi_Deva_kas_Deva 3668955 gom_Deva_san_Deva 942936 kan_Knda_mni_Mtei 2604263 mai_Deva_sat_Olck 360250 mni_Mtei_tel_Telu 8112526
ben_Beng_guj_Gujr 78967455 doi_Deva_mai_Deva 2190770 gom_Deva_sat_Olck 399791 kan_Knda_npi_Deva 3973567 mai_Deva_snd_Arab 1022691 mni_Mtei_urd_Arab 2734445
ben_Beng_hin_Deva 87189520 doi_Deva_mal_Mlym 5447712 gom_Deva_snd_Arab 931740 kan_Knda_ory_Orya 55381962 mai_Deva_snd_Deva 1121301 npi_Deva_ory_Orya 11198582
ben_Beng_kan_Knda 6588212 doi_Deva_mar_Deva 4181626 gom_Deva_snd_Deva 1120965 kan_Knda_pan_Guru 6601794 mai_Deva_tam_Taml 1780322 npi_Deva_pan_Guru 3319821
ben_Beng_kas_Arab 2145259 doi_Deva_mni_Beng 1279927 gom_Deva_tam_Taml 1718214 kan_Knda_san_Deva 3279907 mai_Deva_tel_Telu 11267803 npi_Deva_san_Deva 1565423
ben_Beng_kas_Deva 4232603 doi_Deva_mni_Mtei 2217101 gom_Deva_tel_Telu 13826104 kan_Knda_sat_Olck 928128 mai_Deva_urd_Arab 2769033 npi_Deva_sat_Olck 360263
ben_Beng_mai_Deva 2254028 doi_Deva_npi_Deva 3384768 gom_Deva_urd_Arab 5767926 kan_Knda_snd_Arab 2572544 mal_Mlym_mar_Deva 4080152 npi_Deva_snd_Arab 1382765
ben_Beng_mal_Mlym 5151355 doi_Deva_ory_Orya 26264311 guj_Gujr_hin_Deva 501919187 kan_Knda_snd_Deva 3701804 mal_Mlym_mni_Beng 1005011 npi_Deva_snd_Deva 2383049
ben_Beng_mar_Deva 4356579 doi_Deva_pan_Guru 6035401 guj_Gujr_kan_Knda 45730533 kan_Knda_tam_Taml 4934229 mal_Mlym_mni_Mtei 1994893 npi_Deva_tam_Taml 1780516
ben_Beng_mni_Beng 897884 doi_Deva_san_Deva 2737805 guj_Gujr_kas_Arab 5266708 kan_Knda_tel_Telu 90235065 mal_Mlym_npi_Deva 2336955 npi_Deva_tel_Telu 64765229
ben_Beng_mni_Mtei 2310198 doi_Deva_sat_Olck 849840 guj_Gujr_kas_Deva 10007355 kan_Knda_urd_Arab 8070849 mal_Mlym_ory_Orya 50442690 npi_Deva_urd_Arab 3595423
ben_Beng_npi_Deva 2621108 doi_Deva_snd_Arab 2211042 guj_Gujr_mai_Deva 4852703 kas_Arab_kas_Deva 1799571 mal_Mlym_pan_Guru 5425283 ory_Orya_pan_Guru 48738324
ben_Beng_ory_Orya 52347364 doi_Deva_snd_Deva 2821105 guj_Gujr_mal_Mlym 41472075 kas_Arab_mai_Deva 1140655 mal_Mlym_san_Deva 2397623 ory_Orya_san_Deva 20414334
ben_Beng_pan_Guru 5688148 doi_Deva_tam_Taml 4419242 guj_Gujr_mar_Deva 76034043 kas_Arab_mal_Mlym 2320653 mal_Mlym_sat_Olck 713045 ory_Orya_sat_Olck 3094866
ben_Beng_san_Deva 2876914 doi_Deva_tel_Telu 26541111 guj_Gujr_mni_Beng 1289734 kas_Arab_mar_Deva 2033332 mal_Mlym_snd_Arab 2369219 ory_Orya_snd_Arab 14829766
ben_Beng_sat_Olck 899756 doi_Deva_urd_Arab 6725444 guj_Gujr_mni_Mtei 4373466 kas_Arab_mni_Beng 439988 mal_Mlym_snd_Deva 3280861 ory_Orya_snd_Deva 15923972
ben_Beng_snd_Arab 2322747 eng_Latn_gom_Deva 27718817 guj_Gujr_npi_Deva 30603002 kas_Arab_mni_Mtei 928839 mal_Mlym_tam_Taml 4176661 ory_Orya_tam_Taml 43120335
ben_Beng_snd_Deva 3155804 eng_Latn_guj_Gujr 596740761 guj_Gujr_ory_Orya 217767278 kas_Arab_npi_Deva 1384843 mal_Mlym_tel_Telu 81261194 ory_Orya_tel_Telu 430745670
ben_Beng_tam_Taml 4180758 eng_Latn_hin_Deva 832738952 guj_Gujr_pan_Guru 40878961 kas_Arab_ory_Orya 10958432 mal_Mlym_urd_Arab 6862158 ory_Orya_urd_Arab 61149573
ben_Beng_tel_Telu 87561402 eng_Latn_kan_Knda 115521466 guj_Gujr_san_Deva 14827663 kas_Arab_pan_Guru 2423309 mar_Deva_mni_Beng 869866 pan_Guru_san_Deva 2766113
ben_Beng_urd_Arab 42602867 eng_Latn_kas_Arab 25142330 guj_Gujr_sat_Olck 1541029 kas_Arab_san_Deva 1123349 mar_Deva_mni_Mtei 1763291 pan_Guru_sat_Olck 840630
brx_Deva_doi_Deva 1916922 eng_Latn_kas_Deva 41442708 guj_Gujr_snd_Arab 9319233 kas_Arab_sat_Olck 459615 mar_Deva_npi_Deva 1783813 pan_Guru_snd_Arab 2625050
7 Text Perturbation
We employ various text perturbation techniques to generate task-specific corpora for translation-related applications. By leveraging existing parallel and monolingual corpora, we introduce controlled modifications using carefully designed methods. These denoising strategies, which include random additions, deletions, and substitutions, effectively simulate errors in spelling, grammar, word order, punctuation, and more for all 36 languages mentioned in this work. The crafted perturbation methods are outlined below:
  • Add/Delete/Replace Random Token: In this, random tokens are added, deleted, or replaced to simulate common noise found in real text. This introduces variability and errors in the text. For example, adding irrelevant filler words or dropping crucial terms.
  • Change Pronoun: This perturbation introduces pronoun errors by replacing correct pronouns with incorrect ones, disrupting the coherence and referential clarity of the text. For example, replacing "he" with "she" or "they" can create errors critical for evaluating gender-sensitive, discourse-level translation tasks. To implement this, we used available POS-tagged corpora to compile a list of pronouns for each respective language. For a given input text, a random pronoun is identified and replaced with another randomly selected pronoun from the list.
  • Change Prepositions or Postpositions: By altering prepositions or postpositions, we introduce syntactic inconsistencies and semantic shifts. For example, replacing "on the table" with "under the table" changes the spatial relationship and meaning. To implement this, we used POS-tagged corpora to compile a list of prepositions or postpositions for each language. Similar to the pronoun perturbation, we randomly select a preposition or postposition in the input text and replace it with another from the predefined list. This perturbation is crucial for tasks focusing on syntax and semantic correctness in translation.
  • Change connectives: This perturbation modifies logical connectives, introducing errors that affect the flow and coherence of the text. For instance, replacing "because" with "although" or "and" with "but" creates inconsistencies in the logical structure of sentences. To implement this, we curated a list of commonly used connectives for POS tagged corpora of each language. A random connective in the input text is selected and substituted with another from the list. This method helps with errors in logical and discourse-level translation tasks.
  • Change Verb Forms: This task modifies verb forms to incorrect tense, aspect, or modality. For example, replacing “was running” with “run” or “will ran” introduces errors in temporal and modal constructs that the model must learn to correct.
  • Change Lexical Cohesion: Words or phrases contributing to the overall cohesion of a passage are replaced with less cohesive random alternatives, disrupting the flow and making the sentence appear disjointed. This method helps with errors in logical and discourse-level translation tasks focusing on lexical cohesion.
  • Change Punctuation: Punctuation errors, such as missing periods, misplaced commas, or incorrect quotation marks, are introduced to mimic common typographical errors. Correcting such errors helps improve text readability and grammaticality.
  • Grammar: In this denoising process, selected words are randomly replaced with grammatically inflected forms derived from a trie structure built using language-specific monolingual corpora. This introduces variations in features like gender, number, person, and verb aspects such as tense and modality. The approach is particularly useful for Indian languages with agglutinative properties, enhancing the understanding and modeling of their complex morphological structures to improve language processing and analysis.
  • Masking: In this task, specific words are masked to evaluate the model’s ability to reconstruct missing information. Tokens are randomly replaced with a single random token using a 0.1% probability. Inspired by BART, our masking is applied at the token level before subword tokenization. This approach is particularly challenging for Dravidian languages, requiring the model to predict correct linguistic inflections during decoding.
  • Spelling: Typographical errors are introduced into selected words, affecting around 0.1% of the text. Spelling perturbation involves adding, removing, or substituting random characters, simulating real-world spelling errors. This helps train the model for tasks requiring spelling correction.
  • Word Order: Indian languages often have flexible word order. To address this, we apply word order perturbation by randomly rearranging groups of up to four words with a 0.1% probability. This modifies the original sequence while preserving grammaticality, helping models adapt to structural variations in sentences.

### v2

@@
 BhashaVerse : Translation Ecosystem for Indian Subcontinent Languages
-12024
-[…]
-BhashaVerse : Translation Ecosystem for Indian Subcontinent Languages
+Abstract
+We present a multilingual, multi-task translation model for 36 Indian languages, including Assamese, Hindi, Tamil, and Urdu.
+The model utilizes 10 billion parallel corpora for 36 x 36 language directions and millions of data pairs for tasks like translation evaluation, post-editing, grammar correction, and error identification.
+Corpus creation involves leveraging existing resources, synthetic data, and domain-specific datasets.
+The model’s multi-task approach integrates translation, grammar correction, error identification, quality estimation, and post-editing, ensuring robust cross-lingual communication.
+Evaluations cover discourse-level and domain-specific translation, using reference-based and reference-free metrics.
+This work enhances translation quality, supports linguistic diversity, and advances low-resource language processing.
 issue: 1
@@
 Automatic post-editing systems can then help correct translation errors efficiently, either independently or in collaboration with human translators.
-Incorporating these advancements across general and domain-specific applications will greatly benefit the language research community and end-users.
+Incorporating these advancements across general and domain-specific applications will greatly benefit the language research community and end users.
 This paper focuses on developing translation models and related applications for 36 Indian languages, including Assamese, Awadhi, Bengali, Bhojpuri, Braj, Bodo, Dogri, English, Konkani, Gondi, Gujarati, Hindi, Hinglish, Ho, Kannada, Kangri, Kashmiri (Arabic and Devanagari), Khasi, Mizo, Magahi, Maithili, Malayalam, Marathi, Manipuri (Bengali and Meitei), Nepali, Oriya, Punjabi, Sanskrit, Santali, Sinhala, Sindhi (Arabic and Devanagari), Tamil, Tulu, Telugu, and Urdu.
@@
 By integrating these elements, the study establishes a comprehensive framework to improve machine translation quality and enable better cross-lingual communication in India’s linguistically diverse ecosystem.
-Finally, this work addresses several translation-related applications that are central to enhancing multilingual communication.
+Finally, this work addresses the following translation-related tasks that are central to enhancing multilingual communication in the Indian subcontinent.
   1. 1.
@@
 Our work evaluates how well current systems handle these aspects and proposes improvements for translating connected text effectively.
-Discourse translation possible for 3636 language pairs with keeping 1536 tokens as a context with encoder-decoder model while 4096 context for decoder only model.
+Discourse translation possible for 36 x 36 language pairs with keeping 1536 tokens as a context with encoder-decoder model while 4096 context for decoder only model.
   3. 3.
@@
 We design systems capable of handling domain-specific language nuances to improve usability in practical scenarios.
-We include domains
   4. 4.
@@
 We explore alignment techniques, domain-specific corpora development, synthetic data generation methods, and corpus cleaning methodologies to build more than 1B parallel corpora for machine translation involving 36 Indian subcontinent languages.
-3.1 Automatic Alignment and Human Validation
+Automatic Alignment and Human Validation
 We collected data from various multilingual websites and books containing content in English and multiple Indian languages.
@@
 To address these, we conducted human validation on 10% of the sampled data to assess and adjust the alignment thresholds, ensuring the creation of high-quality parallel corpora suitable for training machine translation systems.
-3.2 Human Post-edited Education and Medical Domain Parallel Corpora
+Human Post-edited Education and Medical Domain Parallel Corpora
 Domain-specific corpora are essential for developing machine translation (MT) systems that perform exceptionally well in specialized areas such as medicine and technical content.
@@
 Notably, the post-edited parallel corpora were developed with English as the source language, and the same English text was post-edited into other Indian languages, enabling the creation of n-way parallel corpora among Indian languages.
-3.3 Quality Synthetic Parallel Corpora
+Quality Synthetic Parallel Corpora
 Synthetic corpora can address the paucity of parallel data for Indian languages.
@@
 These methods significantly enhance the scope of existing corpora, expanding their coverage and offering vital support for low-resource language pairs.
-3.3.1 Pivoted Parallel Corpora
+Pivoted Parallel Corpora
 Most of the parallel corpora developed for Indian languages focus on translations involving English, with relatively few efforts dedicated to creating corpora for Indian-to-Indian language translations.
@@
 This approach facilitates cross-lingual applications and fosters inclusivity in multilingual natural language processing.
-3.3.2 Iterative Backward Translation
+Iterative Backward Translation
 We employed back-translation, a widely used data augmentation technique in machine translation, to enhance the diversity and quality of training data.
@@
 The translated data introduced variations in sentence structure, vocabulary, and phrasing, enriching the training set’s diversity.
-For each monolingual corpus, both translation outputs were evaluated using COMET-QE scores, and the version with the highest score was retained.
-This filtered data was then used to train a multilingual machine translation model involving the selected language pairs.
-The newly trained model was subsequently used to re-translate the same monolingual corpora, iteratively improving the translations.
-This process was repeated for five iterations, with the best-scored translation pairs from all iterations being selected for the final training corpus.
+For each monolingual corpus, both translation outputs were evaluated using COMET-QE scores, and the version with the highest score is retained.
+This filtered data is then used to train a multilingual machine translation model involving the selected language pairs.
+The newly trained model is subsequently used to re-translate the same monolingual corpora, iteratively improving the translations.
+This process is repeated for five iterations, with the best-scored translation pairs from all iterations being selected for the final training corpus.
 Paragraph Level Backward Translation
-For the Wikipedia dataset [^25] https://huggingface.co/datasets/wikimedia/wikipedia and crawled news articles from the internet, we adopted paragraph-level backward translation to preserve coherence and contextual consistency across entire paragraphs.
+For the Wikipedia dataset [^25] https://huggingface.co/datasets/wikimedia/wikipedia and crawled news articles from the internet and from the available monolingual corpora [^26] https://data.statmt.org/news-crawl/, we adopted paragraph-level backward translation to preserve coherence and contextual consistency across entire paragraphs.
 Unlike sentence-level back-translation, this approach ensures that the translated content retains discourse-level attributes, such as topic continuity, pronoun resolution, and lexical cohesion.
 By focusing on paragraph-level fidelity, this method enhances the quality of machine translation systems, particularly for tasks requiring deeper contextual understanding, such as document translation or conversational AI.
-The same filtering mechanism was applied here as well.
-Each sentence within a paragraph was evaluated for quality using COMET-QE scores, and the best-scored translations were selected.
+The same filtering mechanism is applied here as well.
+Each sentence within a paragraph is evaluated for quality using COMET-QE scores, and the best-scored translations were selected.
 These high-quality sentence translations were then reassembled into paragraphs to maintain the original discourse structure.
 This strategy ensured the development of a robust multilingual machine translation model capable of handling complex and contextually rich translations.
-3.4 Parallel Corpora Cleaning
+Parallel Corpora Cleaning
 Cleaning parallel corpora is crucial for ensuring the accuracy and reliability of training datasets.
@@
 COMET-QE Scoring
-The quality of source-target and target-source sentence pairs was evaluated using COMET-22 scores.
-Sentence pairs with scores significantly below the average threshold (average minus 10) were excluded, ensuring that only high-quality data was retained for training.
+The quality of source-target and target-source sentence pairs is evaluated using COMET-22 scores.
+Sentence pairs with scores significantly below the average threshold (average minus 10) are excluded, ensuring that only high-quality data is retained for training.
 We applied these cleaning methods to existing, newly developed, and synthetically generated parallel corpora for Indian languages to achieve greater alignment accuracy, noise reduction, and linguistic consistency, resulting in more effective MT systems.
-4 Automatic Post-Editing (APE) Corpora
-We propose methods to enhance corpora for automatic post-editing between English and Indian languages.
-These include using human post-edited corpora and generating synthetic corpora from existing parallel data, providing a foundation for general and domain-specific post-editing tasks.
-4.1 Human APE Corpora
-As discussed in Section 3.2, post-editing involves human translators refining machine-generated translations to produce accurate and natural-sounding final translations.
-This process was conducted using the PostEditMe [^26] https://posteditme.in/ workbench, a platform designed to facilitate post-editing tasks.
-The workbench allowed translators to edit machine-generated translations while maintaining a record of both the original machine-generated version and the human-corrected version.
-Additionally, the workbench captured all intermediate stages of post-editing, recording multiple iterations for each sentence pair, along with the final edited version.
-These intermediate stages and the final post-edited text form a comprehensive resource for studying the post-editing process and its impact on translation quality.
-The resulting dataset, named the Bhashik APE corpus, includes the initial machine-generated (also intermediate stages text) and human-post-edited translations for source texts in English and 11 Indian languages.
-This corpus supports the development of automatic post-editing systems and enhances domain-specific machine translation by addressing the linguistic nuances of technical and educational contexts more effectively.
-4.2 Synthetically Developed Corpora
-Automatic post-editing (APE) corpora typically consist of three key components: the source text, an imperfect or erroneous translation, and the error-free translation.
-To synthetically generate such APE corpora, we utilized both existing parallel corpora and our newly developed parallel corpora described in this work.
-The target sentences or passages are subjected to various perturbation techniques as detailed in Appendix Section 7, simulating errors commonly found in machine translations.
-For our opted language pairs, we processed each sentence pair to create multiple denoised versions of the target text.
-These perturbations introduced varying degrees of errors, ranging from 2% to 15% of the total words, including grammar mistakes, mistranslations, and word order issues.
-These artificially generated error patterns simulate typical MT errors, making the perturbed version function as the erroneous translation, while the original target text serves as the error-free translation.
-Additionally, we applied noisy back-translation techniques to create diverse post-editing pairs, ensuring a wide range of error types for robust APE training data.
-This approach provides a rich and diverse dataset for training automatic post-editing models.
-We call the resulting dataset Bhashik APE (Synth).
-5 Corpora for Machine Translation Evaluation
-In this work, we developed machine translation assessment corpora for English and Indian languages through multiple methods to evaluate translation quality.
-We focus on translation evaluation which is on a continuous scale from 1 to 100, where 1 indicates poor translation and 100 represents the best possible translation.
-We use two approaches: reference-based evaluation, comparing translations to a reference, and reference-free evaluation, assessing translations without references.
-Corpora development involves human expert scoring and synthetic data generation using perturbation techniques (Appendix Section 7) to simulate common translation errors.
-Human assessment corpora are created by experts scoring translations on a continuous scale, while synthetic data is generated with COMET models scoring translations across scenarios derived from parallel corpora.
-This section outlines the strategies for creating synthetic corpora to comprehensively evaluate machine translation systems.
-5.1 Human Annotated Corpora
-Human-annotated corpora are created by language experts who evaluate translations based on fluency, adequacy, and contextual accuracy, providing a gold standard for assessing MT systems.
-Experts fluent in both the source and target languages assess the accuracy of the translation in conveying the original meaning, focusing on Translation Accuracy, which determines whether the translation preserves the source text’s meaning while maintaining fluency and grammatical correctness.
-Scores are assigned on a scale of 1 to 100.
-To develop these corpora, we utilized benchmark datasets and multiple translation engines, including Google Translate, SSMTv3, and Indictransv2, to generate translations for input texts.
-Experts evaluated both the source text and the translated text, providing ratings.
-The reference translations in the benchmark corpora were used for reference-enabled machine translation evaluation, while the annotations can also be applied in reference-free evaluation scenarios.
-The dataset spans 18 language pairs and 36 translation directions, resulting in over 72,000 rated translations.
-This effort involved over 70 experienced language experts, with each source and target translation pair receiving three independent ratings.
-5.2 Synthetically Developed Corpora
-To achieve comprehensive coverage of the translation evaluation spectrum on a linear scale, we recognized that relying solely on existing or limited human scored direct assessment data is insufficient.
-Therefore, we decided to generate additional data synthetically using existing parallel and benchmark corpora.
-Various techniques are employed to create degraded, perturbed versions of translations, and we used an existing COMET model to assign degraded scores using the following methodology:
-  1. 1.
-    Generating Perturbed Translations: We began by sourcing parallel and benchmark corpora from above mentioned repositories.
-For each sentence pair, we applied perturbation techniques outlined in Appendix Section 7 to generate lower-quality translations with varying degrees of errors.
-The error levels ranged from 5% to 50%.
-  2. 2.
-    Structuring Sentence Tuples: For each perturbed version, we created a sentence tuple containing the following elements:
-    • Source text (S)
-    • Original translation text (T) from the parallel corpora
-    • Perturbed translation text (T̃) with Error percentage (Ep)
-  3. 3.
-    Calculating Synthetic Scores: Using the existing COMET Quality Estimation (QE) model, we first evaluated the source text (S) and the original translation (T) to obtain a COMET score (Cs) and 100 minus the Translation Error Rate (TER) score (Tp) using original translation (T) and Perturbed translation text (T̃).
-We then averaged the error percentage (Ep) and the TER-derived score (Tp) to calculate a degradation factor.
-This factor was subtracted from the COMET score (Cs) to derive a new synthetic score for the source text (S) and the perturbed translation (T̃).
-By applying this process, we generated an extensive synthetic dataset to supplement the limited human-annotated data, enabling robust evaluation across a broader range of translation quality scenarios.
-The resulting dataset is named the Bhashik PseudoMTDA corpus for English and 11 Indian languages.
-6 Corpora for Machine Translation Error Span Identification and Error Category Marking
-Error identification and categorization help analyze weaknesses in MT systems, enabling targeted improvements. This section outlines our development of corpora to support these tasks.
-6.1 Human MT Error Span Identification Corpora
-As discussed in Section 3.2, post-editing involves human translators refining machine-generated translations to produce accurate, natural-sounding final translations.
-This process was carried out using the PostEditMe [^27] https://posteditme.in/ workbench, a platform specifically designed to support post-editing tasks.
-The workbench captures every edit made by translation experts and records this information at the word level over the machine translation output.
-The edited spans in the translated text, where changes were made, are marked as MT error spans for the corresponding translation.
-These post-editing efforts provide a rich resource for identifying error spans and evaluating translation quality.
-The resulting dataset, named the Bhashik PseudoMTError corpus, includes the source text, the initial machine-generated translation, and the annotated error spans for English and 11 Indian languages.
-6.2 Synthetically Developed MT Error Span and Category Marking Corpora
-As discussed earlier, corpora for machine translation (MT) error span and category marking typically include three key components: the source text, a translation annotated with error spans, and corresponding error categories.
-To synthetically create such corpora, we leveraged both existing parallel corpora and newly developed parallel corpora introduced in this work.
-Target sentences or passages were subjected to various perturbation techniques, detailed in Appendix Section 7, to simulate common machine translation errors across different categories.
-These categories were aligned with broader MQM categories.
-For the selected language pairs, we processed each sentence pair to generate multiple denoised versions of the target text.
-Perturbations introduced errors ranging from 2% to 15% of the total words, including grammar mistakes, mistranslations, and word order issues.
-These artificially generated error patterns mimic typical MT errors, making the perturbed versions serve as erroneous translations.
-Error spans were annotated based on the introduced errors, along with the corresponding categories used to generate the perturbations.
-This approach provides a rich and diverse dataset for training automatic systems to mark MT error spans and categories. We refer to the resulting dataset as the Bhashik PseudoMTError Corpus.
 3.1 Developed Parallel Corpora
-6.3.1 Bhashik Education and Medical: Human Post-edited Corpora
-Table 1: The post-edited 2M domain-parallel corpora spans English and 10 Indian languages, covering 15 educational domains.
-Domain Language Pair Passages/Sentences Domain Language Pair Passages/Sentences
-Natural science eng_Latn-hin_Deva 140999 Teaching methods eng_Latn-tam_Taml 17
-Natural science eng_Latn-mar_Deva 191936 Teaching methods eng_Latn-mal_Mlym 527
-Natural science eng_Latn-kan_Knda 185491 Teaching methods eng_Latn-urd_Arab 3909
-Natural science eng_Latn-tel_Telu 223146 Management eng_Latn-mar_Deva 14521
-Natural science eng_Latn-mal_Mlym 84 Management eng_Latn-kan_Knda 11001
-Natural science eng_Latn-tam_Taml 38 Management eng_Latn-tel_Telu 11041
-Computer science eng_Latn-mar_Deva 80170 Health eng_Latn-hin_Deva 715
-Computer science eng_Latn-hin_Deva 71507 Psychology eng_Latn-hin_Deva 12651
-Computer science eng_Latn-kan_Knda 75354 Psychology eng_Latn-tel_Telu 12771
-Computer science eng_Latn-tel_Telu 75388 Psychology eng_Latn-mar_Deva 12822
-Computer science eng_Latn-urd_Arab 332 Psychology eng_Latn-kan_Knda 12570
-Computer science eng_Latn-tam_Taml 11902 Marketing eng_Latn-hin_Deva 4903
-Law eng_Latn-tel_Telu 80165 Marketing eng_Latn-tel_Telu 4971
-Law eng_Latn-mar_Deva 81581 Marketing eng_Latn-mar_Deva 4858
-Law eng_Latn-tam_Taml 90 Marketing eng_Latn-kan_Knda 5099
-Law eng_Latn-kan_Knda 78245 Political science eng_Latn-tel_Telu 11978
-Law eng_Latn-hin_Deva 735 Political science eng_Latn-mar_Deva 11935
-Mathematics eng_Latn-mar_Deva 62881 Political science eng_Latn-kan_Knda 12001
-Mathematics eng_Latn-tel_Telu 59373 Political science eng_Latn-hin_Deva 7876
-Mathematics eng_Latn-kan_Knda 40625 Textile eng_Latn-tel_Telu 7609
-Mathematics eng_Latn-hin_Deva 26259 Textile eng_Latn-mar_Deva 7637
-Mathematics eng_Latn-tam_Taml 286 Textile eng_Latn-kan_Knda 7630
-Mathematics eng_Latn-mal_Mlym 163 History eng_Latn-tel_Telu 4902
-health eng_Latn-tam_Taml 12728 History eng_Latn-mar_Deva 4919
-health eng_Latn-guj_Gujr 299 History eng_Latn-kan_Knda 4941
-health eng_Latn-tel_Telu 34855 Economics eng_Latn-tel_Telu 5087
-health eng_Latn-hin_Deva 11802 Economics eng_Latn-mar_Deva 5093
-health eng_Latn-mar_Deva 1904 Economics eng_Latn-kan_Knda 5061
-health eng_Latn-kan_Knda 135 Tourism eng_Latn-tel_Telu 12798
-health eng_Latn-mal_Mlym 202 Tourism eng_Latn-mar_Deva 12795
-health eng_Latn-ben_Beng 370 Tourism eng_Latn-kan_Knda 12813
-Teaching methods eng_Latn-kan_Knda 32670 Communication skill eng_Latn-tel_Telu 8426
-Teaching methods eng_Latn-tel_Telu 31828 Communication skill eng_Latn-mar_Deva 8537
-Teaching methods eng_Latn-hin_Deva 19065 Communication skill eng_Latn-kan_Knda 8627
-Teaching methods eng_Latn-mar_Deva 32327
-6.3.2 Bhashik Generic : Quality Synthetic Corpora
-Table 2: The generated 1B parallel corpora cover 325 language pairs, encompassing 36 languages from the Indian subcontinent and utilizing 15 unique scripts.
+Table 1: The generated over 1B parallel corpora cover 325 language pairs, encompassing 36 languages from the Indian subcontinent and utilizing 15 unique scripts.
 Language Pair
@@
 brx_Deva_doi_Deva 1916922 eng_Latn_kas_Deva 41442708 guj_Gujr_snd_Arab 9319233 kas_Arab_sat_Olck 459615 mar_Deva_npi_Deva 1783813 pan_Guru_snd_Arab 2625050
-7 Text Perturbation
-We employ various text perturbation techniques to generate task-specific corpora for translation-related applications.
-By leveraging existing parallel and monolingual corpora, we introduce controlled modifications using carefully designed methods.
-These denoising strategies, which include random additions, deletions, and substitutions, effectively simulate errors in spelling, grammar, word order, punctuation, and more for all 36 languages mentioned in this work.
-The crafted perturbation methods are outlined below:
+Table 2: Postedited Health Domain Corpora for English and 8 Indian Languages
+Language Pair Passages/Sentences Language Pair Passages/Sentences
+ben_Beng_Latn-tam_Taml 142 eng_Latn-tel_Telu 10308
+ben_Beng_Latn-tel_Telu 291 hin_Deva-ben_Beng_Latn 6384
+eng_Latn-ben_Beng_Latn 7915 hin_Deva-guj_Gujr 145
+eng_Latn-guj_Gujr 158 hin_Deva-kan_Knda 1138
+eng_Latn-hin_Deva 11042 hin_Deva-mal_Mlym 1436
+eng_Latn-kan_Knda 1209 hin_Deva-mar_Deva 879
+eng_Latn-mal_Mlym 1754 hin_Deva-tam_Taml 5517
+eng_Latn-mar_Deva 906 hin_Deva-tel_Telu 6412
+eng_Latn-tam_Taml 8726 tam_Taml-tel_Telu 1019
+Table 3: The human post-edited 2M domain-parallel corpora spans English and 10 Indian languages, covering 15 educational domains.
+Domain Language Pair Passages/Sentences Domain Language Pair Passages/Sentences
+Natural science eng_Latn-hin_Deva 140999 Teaching methods eng_Latn-tam_Taml 17
+Natural science eng_Latn-mar_Deva 191936 Teaching methods eng_Latn-mal_Mlym 527
+Natural science eng_Latn-kan_Knda 185491 Teaching methods eng_Latn-urd_Arab 3909
+Natural science eng_Latn-tel_Telu 223146 Management eng_Latn-mar_Deva 14521
+Natural science eng_Latn-mal_Mlym 84 Management eng_Latn-kan_Knda 11001
+Natural science eng_Latn-tam_Taml 38 Management eng_Latn-tel_Telu 11041
+Computer science eng_Latn-mar_Deva 80170 Health eng_Latn-hin_Deva 715
+Computer science eng_Latn-hin_Deva 71507 Psychology eng_Latn-hin_Deva 12651
+Computer science eng_Latn-kan_Knda 75354 Psychology eng_Latn-tel_Telu 12771
+Computer science eng_Latn-tel_Telu 75388 Psychology eng_Latn-mar_Deva 12822
+Computer science eng_Latn-urd_Arab 332 Psychology eng_Latn-kan_Knda 12570
+Computer science eng_Latn-tam_Taml 11902 Marketing eng_Latn-hin_Deva 4903
+Law eng_Latn-tel_Telu 80165 Marketing eng_Latn-tel_Telu 4971
+Law eng_Latn-mar_Deva 81581 Marketing eng_Latn-mar_Deva 4858
+Law eng_Latn-tam_Taml 90 Marketing eng_Latn-kan_Knda 5099
+Law eng_Latn-kan_Knda 78245 Political science eng_Latn-tel_Telu 11978
+Law eng_Latn-hin_Deva 735 Political science eng_Latn-mar_Deva 11935
+Mathematics eng_Latn-mar_Deva 62881 Political science eng_Latn-kan_Knda 12001
+Mathematics eng_Latn-tel_Telu 59373 Political science eng_Latn-hin_Deva 7876
+Mathematics eng_Latn-kan_Knda 40625 Textile eng_Latn-tel_Telu 7609
+Mathematics eng_Latn-hin_Deva 26259 Textile eng_Latn-mar_Deva 7637
+Mathematics eng_Latn-tam_Taml 286 Textile eng_Latn-kan_Knda 7630
+Mathematics eng_Latn-mal_Mlym 163 History eng_Latn-tel_Telu 4902
+health eng_Latn-tam_Taml 12728 History eng_Latn-mar_Deva 4919
+health eng_Latn-guj_Gujr 299 History eng_Latn-kan_Knda 4941
+health eng_Latn-tel_Telu 34855 Economics eng_Latn-tel_Telu 5087
+health eng_Latn-hin_Deva 11802 Economics eng_Latn-mar_Deva 5093
+health eng_Latn-mar_Deva 1904 Economics eng_Latn-kan_Knda 5061
+health eng_Latn-kan_Knda 135 Tourism eng_Latn-tel_Telu 12798
+health eng_Latn-mal_Mlym 202 Tourism eng_Latn-mar_Deva 12795
+health eng_Latn-ben_Beng 370 Tourism eng_Latn-kan_Knda 12813
+Teaching methods eng_Latn-kan_Knda 32670 Communication skill eng_Latn-tel_Telu 8426
+Teaching methods eng_Latn-tel_Telu 31828 Communication skill eng_Latn-mar_Deva 8537
+Teaching methods eng_Latn-hin_Deva 19065 Communication skill eng_Latn-kan_Knda 8627
+Teaching methods eng_Latn-mar_Deva 32327
+In this section, we introduce the corpora we developed for machine translation between English and 35 Indian languages, covering both general and domain-specific contexts, as detailed in Tables 1, 3 and 2.
+These parallel corpora are publicly available for research purposes on Hugging Face  [^27] https://huggingface.co/ltrciiith under the names: bhashik-parallel-corpora-generic, bhashik-parallel-corpora-education, and bhashik-parallel-corpora-health.
+4 Automatic Post-Editing (APE) Corpora
+We propose methods to enhance corpora for automatic post-editing between English and Indian languages.
+These include using human post-edited corpora and generating synthetic corpora from existing parallel data, providing a foundation for general and domain-specific post-editing tasks.
+Table 4: Automatic post-editing corpora for English and Indian languages: general domain (Bhashik APE Synth - 10M) and education and health domains (Bhashik APE Human - 2M).
+Bhashik APE (Synth) : General Bhashik APE (Synth) : General Bhashik APE (Human) : Education and Health
+Source Language Target Language #Sentences Source Language Target Language #Sentences Source Language Target Language #Sentences
+eng_Latn san_Deva 167912 asm_Beng eng_Latn 307940 eng_Latn tam_Taml 13958
+eng_Latn mni_Mtei 75170 ben_Beng eng_Latn 498159 eng_Latn mar_Deva 733974
+eng_Latn ory_Orya 188394 brx_Deva eng_Latn 221311 eng_Latn kan_Knda 664516
+eng_Latn doi_Deva 95643 doi_Deva eng_Latn 197047 eng_Latn hin_Deva 407093
+eng_Latn mai_Deva 124677 gom_Deva eng_Latn 201763 eng_Latn urd_Arab 4307
+eng_Latn tam_Taml 113849 guj_Gujr eng_Latn 327686 eng_Latn tel_Telu 751077
+eng_Latn mar_Deva 297651 hin_Deva eng_Latn 564166 eng_Latn mal_Mlym 999
+eng_Latn kan_Knda 135478 kan_Knda eng_Latn 432813
+eng_Latn npi_Deva 231413 kas_Arab eng_Latn 196620
+eng_Latn hin_Deva 261516 kas_Deva eng_Latn 43189
+eng_Latn urd_Arab 225262 mai_Deva eng_Latn 156935
+eng_Latn asm_Beng 258234 mal_Mlym eng_Latn 524830
+eng_Latn guj_Gujr 146769 mar_Deva eng_Latn 274688
+eng_Latn brx_Deva 126492 mni_Beng eng_Latn 42669
+eng_Latn ben_Beng 264087 mni_Mtei eng_Latn 203056
+eng_Latn gom_Deva 88346 npi_Deva eng_Latn 561283
+eng_Latn tel_Telu 192549 ory_Orya eng_Latn 385591
+eng_Latn sat_Olck 43932 pan_Guru eng_Latn 293203
+eng_Latn kas_Arab 45590 san_Deva eng_Latn 217712
+eng_Latn mal_Mlym 723 sat_Olck eng_Latn 139586
+tel_Telu eng_Latn 465576 snd_Arab eng_Latn 24459
+urd_Arab eng_Latn 457598 snd_Deva eng_Latn 93589
+tam_Taml eng_Latn 312192
+Human APE Corpora
+As discussed in Section 3, post-editing involves human translators refining machine-generated translations to produce accurate and natural-sounding final translations.
+This task was conducted using the PostEditMe [^28] https://posteditme.in/ workbench, a platform designed to facilitate post-editing tasks.
+The workbench allowed translators to edit machine-generated translations while maintaining a record of both the original machine-generated version and the human-corrected version.
+Additionally, the workbench captured all intermediate stages of post-editing, recording multiple iterations for each sentence pair, along with the final edited version.
+These intermediate stages and the final post-edited text form a comprehensive resource for studying the post-editing process and its impact on translation quality.
+The resulting dataset, named the Bhashik APE corpus (presented in Table 4), includes the initial machine-generated (also intermediate stages text) and human-post-edited translations for source texts in English and 11 Indian languages.
+This corpus supports the development of automatic post-editing systems and enhances domain-specific machine translation by addressing the linguistic nuances of technical and educational contexts more effectively.
+Synthetically Developed Corpora
+Automatic post-editing (APE) corpora typically consist of three key components: the source text, an imperfect or erroneous translation, and the error-free translation.
+To synthetically generate such APE corpora, we utilized both existing parallel corpora and our newly developed parallel corpora described in this work.
+The target sentences or passages are subjected to various perturbation techniques as detailed in Appendix Section 7, simulating errors commonly found in machine translations.
+For our opted language pairs, we processed each sentence pair to create multiple denoised versions of the target text.
+These perturbations introduced varying degrees of errors, ranging from 2% to 15% of the total words, including grammar mistakes, mistranslations, and word order issues.
+These artificially generated error patterns simulate typical MT errors, making the perturbed version function as the erroneous translation, while the original target text serves as the error-free translation.
+Additionally, we applied noisy back-translation techniques to create diverse post-editing pairs, ensuring a wide range of error types for robust APE training data.
+This approach provides a rich and diverse dataset for training automatic post-editing models.
+We call the resulting dataset Bhashik APE (Synth) and details are presented in Table 4.
+5 Corpora for Machine Translation Evaluation
+In this work, we developed machine translation assessment corpora for English and Indian languages through multiple methods to evaluate translation quality.
+We focus on translation evaluation which is on a continuous scale from 1 to 100, where 1 indicates poor translation and 100 represents the best possible translation.
+We use two approaches: reference-based evaluation, comparing translations to a reference, and reference-free evaluation, assessing translations without references.
+Corpora development involves human expert scoring and synthetic data generation using perturbation techniques (Appendix Section 7) to simulate common translation errors.
+Human assessment corpora are created by experts scoring translations on a continuous scale, while synthetic data is generated with COMET models scoring translations across scenarios derived from parallel corpora.
+This section outlines the strategies for creating synthetic corpora to comprehensively evaluate machine translation systems.
+Table 5: 1M Human direct assessment corpora (without reference) for English and Indian languages across general domain for 36 language pairs.
+Source Language Target Language #Sentences Source Language Target Language #Sentences
+eng_Latn ben_Beng 2960 hin_Deva ben_Beng 3001
+eng_Latn guj_Gujr 2961 hin_Deva guj_Gujr 2961
+eng_Latn hin_Deva 2961 hin_Deva kan_Knda 3001
+eng_Latn kan_Knda 3001 hin_Deva mar_Deva 3201
+eng_Latn mar_Deva 3001 hin_Deva ory_Orya 2399
+eng_Latn ory_Orya 2361 hin_Deva pan_Guru 3000
+eng_Latn Panjabi 2961 hin_Deva tel_Telu 3041
+eng_Latn tam_Taml 2961 hin_Deva urd_Arab 2053
+eng_Latn tel_Telu 2961 ben_Beng hin_Deva 3001
+eng_Latn urd_Arab 1994 guj_Gujr hin_Deva 2961
+ben_Beng eng_Latn 2960 kan_Knda hin_Deva 3001
+guj_Gujr eng_Latn 2961 mar_Deva hin_Deva 3201
+hin_Deva eng_Latn 2961 ory_Orya hin_Deva 2080
+kan_Knda eng_Latn 3001 pan_Guru hin_Deva 3000
+mar_Deva eng_Latn 3001 tel_Telu hin_Deva 3041
+ory_Orya eng_Latn 2321 urd_Arab hin_Deva 2001
+pan_Guru eng_Latn 2961 tel_Telu eng_Latn 2961
+tam_Taml eng_Latn 2961 urd_Arab eng_Latn 1961
+Human Annotated Corpora
+Human-annotated corpora are created by language experts who evaluate translations based on fluency, adequacy, and contextual accuracy, providing a gold standard for assessing MT systems.
+Experts fluent in both the source and target languages assess the accuracy of the translation in conveying the original meaning, focusing on Translation Accuracy, which determines whether the translation preserves the source text’s meaning while maintaining fluency and grammatical correctness.
+Scores are assigned on a scale of 1 to 100.
+To develop these corpora, we utilized benchmark datasets and multiple translation engines, including Google Translate, SSMTv3, and Indictransv2, to generate translations for input texts.
+Experts evaluated both the source text and the translated text, providing ratings.
+The reference translations in the benchmark corpora were used for reference-enabled machine translation evaluation, while the annotations can also be applied in reference-free evaluation scenarios.
+The dataset spans 18 language pairs and 36 translation directions, resulting in over 101K rated translations.
+This effort involved over 70 experienced language experts, with each source and target translation pair receiving three independent ratings.
+The human-annotated corpora details are presented in Table 5.
+Table 6: 10M pseudo direct assessment corpora (with reference) for English and Indian languages across general, educational, and health domains.
+Source Language Target Language #Sentences Source Language Target Language #Sentences
+asm_Beng eng_Latn 275688 eng_Latn hin_Deva 668620
+ben_Beng eng_Latn 298946 eng_Latn mar_Deva 1031650
+brx_Deva eng_Latn 139431 eng_Latn tel_Telu 943632
+doi_Deva eng_Latn 113633 eng_Latn urd_Arab 229569
+gom_Deva eng_Latn 112863 eng_Latn kan_Knda 800024
+guj_Gujr eng_Latn 154981 eng_Latn tam_Taml 127947
+hin_Deva eng_Latn 356329 eng_Latn ben_Beng 264087
+kan_Knda eng_Latn 354733 eng_Latn asm_Beng 258234
+kas_Arab eng_Latn 138187 eng_Latn ory_Orya 188394
+kas_Deva eng_Latn 42465 eng_Latn mai_Deva 124677
+mai_Deva eng_Latn 148104 eng_Latn brx_Deva 126492
+mal_Mlym eng_Latn 258292 eng_Latn doi_Deva 95643
+mar_Deva eng_Latn 159852 eng_Latn npi_Deva 231413
+mni_Beng eng_Latn 42324 eng_Latn guj_Gujr 146769
+mni_Mtei eng_Latn 121557 eng_Latn mni_Mtei 75170
+npi_Deva eng_Latn 285402 eng_Latn san_Deva 167912
+ory_Orya eng_Latn 206850 eng_Latn gom_Deva 88346
+pan_Guru eng_Latn 37968 eng_Latn kas_Arab 45590
+san_Deva eng_Latn 171657 eng_Latn sat_Olck 43932
+sat_Olck eng_Latn 139451 eng_Latn mal_Mlym 1722
+snd_Deva eng_Latn 62937 tel_Telu eng_Latn 331194
+tam_Taml eng_Latn 129059 urd_Arab eng_Latn 255991
+Table 7: 70M pseudo direct assessment corpora (without reference) for English and Indian languages across general domain.
+Source Language Target Language #Sentences Source Language Target Language #Sentences
+asm_Beng eng_Latn 615562 eng_Latn ben_Beng 3050599
+ben_Beng eng_Latn 101559 eng_Latn hin_Deva 4949515
+bho_Deva eng_Latn 2991 eng_Latn guj_Gujr 316079
+brx_Deva eng_Latn 25297 eng_Latn mar_Deva 214706
+doi_Deva eng_Latn 115211 eng_Latn pan_Guru 13973
+gom_Deva eng_Latn 23858 eng_Latn kan_Knda 56693
+guj_Gujr eng_Latn 307059 eng_Latn tam_Taml 10747282
+hin_Deva eng_Latn 11386159 eng_Latn tel_Telu 307867
+kan_Knda eng_Latn 63748 eng_Latn mag_Deva $92$
+kas_Arab eng_Latn 149232 eng_Latn mal_Mlym 8963878
+kas_Deva eng_Latn 202814 eng_Latn npi_Deva 51496
+lus_Latn eng_Latn 3021 eng_Latn kas_Deva 33387
+mai_Deva eng_Latn 92319 eng_Latn brx_Deva 50430
+mal_Mlym eng_Latn 8883930 eng_Latn san_Deva 277518
+mar_Deva eng_Latn 75266 eng_Latn kas_Arab 149251
+mni_Beng eng_Latn 8775 eng_Latn gom_Deva 27831
+mni_Mtei eng_Latn 25508 eng_Latn sat_Olck 25126
+npi_Deva eng_Latn 1635019 eng_Latn mni_Mtei 25471
+ory_Orya eng_Latn 2959267 eng_Latn asm_Beng 109159
+pan_Guru eng_Latn 27787 eng_Latn bho_Deva 2955
+san_Deva eng_Latn 277532 eng_Latn ory_Orya 2936801
+sat_Olck eng_Latn 25140 eng_Latn lus_Latn 3006
+sin_Sinh eng_Latn 3045 eng_Latn mai_Deva 30065
+snd_Arab eng_Latn 5688 eng_Latn sin_Sinh 3012
+snd_Deva eng_Latn 13114 eng_Latn mni_Beng 8855
+tam_Taml eng_Latn 10614897 eng_Latn snd_Arab 5575
+tel_Telu eng_Latn 351303 eng_Latn urd_Arab 46907
+urd_Arab eng_Latn 47896 eng_Latn doi_Deva 24177
+eng_Latn hoc_Wara 201 eng_Latn snd_Deva 13027
+Synthetically Developed Corpora
+To achieve comprehensive coverage of the translation evaluation spectrum on a linear scale, we recognized that relying solely on existing or limited human scored direct assessment data is insufficient.
+Therefore, we decided to generate additional data synthetically using existing parallel and benchmark corpora.
+Various techniques are employed to create degraded, perturbed versions of translations, and we used an existing COMET model to assign degraded scores using the following methodology:
+  1. 1.
+    Generating Perturbed Translations: We began by sourcing parallel and benchmark corpora from above mentioned repositories.
+For each sentence pair, we applied perturbation techniques outlined in Appendix Section 7 to generate lower-quality translations with varying degrees of errors.
+The error levels ranged from 5% to 50%.
+  2. 2.
+    Structuring Sentence Tuples: For each perturbed version, we created a sentence tuple containing the following elements:
+    • Source text (S)
+    • Original translation text (T) from the parallel corpora
+    • Perturbed translation text (T̃) with Error percentage (Ep)
+  3. 3.
+    Calculating Synthetic Scores: Using the existing COMET Quality Estimation (QE) model, we first evaluated the source text (S) and the original translation (T) to obtain a COMET score (Cs) and 100 minus the Translation Error Rate (TER) score (Tp) using original translation (T) and Perturbed translation text (T̃).
+We then averaged the error percentage (Ep) and the TER-derived score (Tp) to calculate a degradation factor.
+This factor is subtracted from the COMET score (Cs) to derive a new synthetic score for the source text (S) and the perturbed translation (T̃).
+By applying this process, we generated an extensive synthetic dataset to supplement the limited human-annotated data, enabling robust evaluation across a broader range of translation quality scenarios.
+The resulting dataset is named the Bhashik PseudoMTDA corpus for English and 11 Indian languages.
+The developed corpora for direct assessment, both with and without reference, are shown in Table 6 and Table 7, respectively.
+6 Corpora for MT Error Span Identification
+Table 8: 20M pseudo error-annotated corpora with 54M error spans for English and Indian languages across general, educational, and health domains.
+Source Language Target Language #Sentences #Errors #Errors per Sent Source Language Target Language #Sentences #Errors #Errors per Sent
+asm_Beng eng_Latn 551376 1278030 2.32 eng_Latn hin_Deva 1337218 4849894 3.63
+ben_Beng eng_Latn 597892 1385958 2.32 eng_Latn ben_Beng 528174 1205310 2.28
+brx_Deva eng_Latn 278862 633800 2.27 eng_Latn mar_Deva 2063250 8725642 4.23
+doi_Deva eng_Latn 227266 508742 2.24 eng_Latn npi_Deva 727670 926702 1.27
+gom_Deva eng_Latn 225726 493676 2.19 eng_Latn asm_Beng 516468 1237170 2.4
+guj_Gujr eng_Latn 309962 705820 2.28 eng_Latn san_Deva 337358 1875910 5.56
+hin_Deva eng_Latn 712636 1225358 1.72 eng_Latn mni_Mtei 150340 479422 3.19
+kan_Knda eng_Latn 707896 1022196 1.44 eng_Latn urd_Arab 459138 816608 1.78
+kas_Arab eng_Latn 276374 663646 2.4 eng_Latn mai_Deva 249354 597684 2.4
+kas_Deva eng_Latn 84930 251870 2.97 eng_Latn brx_Deva 252984 684588 2.71
+mai_Deva eng_Latn 296208 665144 2.25 eng_Latn tel_Telu 1887252 6017654 3.19
+mal_Mlym eng_Latn 516584 1189960 2.3 eng_Latn kas_Arab 141542 244316 1.73
+mni_Beng eng_Latn 84648 251446 2.97 eng_Latn ory_Orya 379626 869170 2.29
+mni_Mtei eng_Latn 243114 539906 2.22 eng_Latn tam_Taml 270474 375414 1.39
+npi_Deva eng_Latn 570804 1317958 2.31 eng_Latn doi_Deva 191286 481650 2.52
+ory_Orya eng_Latn 413792 942118 2.28 eng_Latn guj_Gujr 293538 620304 2.11
+pan_Guru eng_Latn 75936 158936 2.09 eng_Latn kan_Knda 1599988 5099336 3.19
+san_Deva eng_Latn 343314 774460 2.26 eng_Latn gom_Deva 353456 296182 0.84
+sat_Olck eng_Latn 278902 620176 2.22 eng_Latn sat_Olck 87864 201198 2.29
+snd_Deva eng_Latn 125874 277914 2.21 eng_Latn mal_Mlym 311942 3244 0.01
+tam_Taml eng_Latn 258118 584962 2.27 tel_Telu eng_Latn 661054 954970 1.44
+urd_Arab eng_Latn 511982 1183818 2.31
+Error identification and categorization help analyze weaknesses in MT systems, enabling targeted improvements.
+This section outlines our development of corpora to support these tasks.
+As discussed in Section 3, post-editing involves human translators refining machine-generated translations to produce accurate, natural-sounding final translations.
+This task was carried out using the PostEditMe [^29] https://posteditme.in/ workbench, a platform specifically designed to support post-editing tasks.
+The workbench captures every edit made by translation experts and records this information at the word level over the machine translation output.
+The edited spans in the translated text, where changes were made, are marked as MT error spans for the corresponding translation.
+These post-editing efforts provide a rich resource for identifying error spans and evaluating translation quality.
+The resulting dataset, named the Bhashik PseudoMTError corpus, includes the source text, the initial machine-generated translation, and the annotated error spans for English and multiple Indian languages.
+Synthetically Developed MT Error Span Marked Corpora
+As discussed earlier, corpora for machine translation (MT) error span and category marking typically include three key components: the source text, a translation annotated with error spans, and corresponding error categories.
+To synthetically create such corpora, we leveraged both existing parallel corpora and newly developed parallel corpora introduced in this work.
+Target sentences or passages were subjected to various perturbation techniques, detailed in Appendix Section 7, to simulate common machine translation errors across different categories.
+These categories were aligned with broader MQM categories.
+For the selected language pairs, we processed each sentence pair to generate multiple denoised versions of the target text.
+Perturbations introduced errors ranging from 2% to 15% of the total words, including grammar mistakes, mistranslations, and word order issues.
+These artificially generated error patterns mimic typical MT errors, making the perturbed versions serve as erroneous translations.
+Error spans were annotated based on the introduced errors to generate the perturbations.
+This approach provides a rich and diverse dataset for training automatic systems to mark MT error spans and categories.
+We refer to the resulting dataset as the Bhashik PseudoMTError Corpus and details are presented in Table 8.
+7 Corpora for Grammar Correction
+We use text perturbation techniques to generate grammar correction corpora by introducing controlled errors in monolingual data.
+These include spelling, grammar, word order, and punctuation errors for all 36 languages.
+The methods are outlined below:
   • Add/Delete/Replace Random Token: In this, random tokens are added, deleted, or replaced to simulate common noise found in real text.
@@
 This modifies the original sequence while preserving grammaticality, helping models adapt to structural variations in sentences.
+8 Languages
+India’s linguistic diversity is rooted in the historical relationships and evolution of languages, shaped by their scripts, families, and groups [^30] https://language.census.gov.in/.
+Understanding these relationships is essential for developing multilingual models and translation systems.
+To address this, we modeled 36 languages using their native scripts, ensuring accurate representation and processing as follows:
+8.1 Language Script
+Scripts visually represent languages, influencing orthography and phonetics.
+Shared scripts, like Devanagari [^31] https://en.wikipedia.org/wiki/Devanagari (used by Hindi, Marathi, Maithili, Bhojpuri, etc.), simplify subword tokenization and multilingual representation.
+Similarly, the Bengali [^32] https://en.wikipedia.org/wiki/Bengali˙alphabet script (used by Bengali, Assamese, Manipuri, etc.) facilitates shared processing.
+However, languages with distinct scripts, such as Tamil [^33] https://en.wikipedia.org/wiki/Tamil˙language (Tamil script), Kannada [^34] https://en.wikipedia.org/?title=Kannada (Kannada script), and Malayalam [^35] https://en.wikipedia.org/wiki/Malayalam (Malayalam script), require different subword vocabularies due to their unique orthographic structures.
+Additionally, languages like Santali [^36] https://en.wikipedia.org/wiki/Santali˙language (Ol Chiki script) and Ho [^37] https://en.wikipedia.org/wiki/Ho˙language (Warang Citi script) reflect unique cultural identities, necessitating custom handling.
+To enable interoperable multilingual systems, script-specific features are mapped to unified representations, balancing shared and unique linguistic properties as part of a custom subword-based tokenizer.
+Table 9: Language Scripts, IDs, and Groupings by Language Family
+ID Language Code Language Group Language Name (script)
+1 asm_Beng Magadhi Assamese (Bengali script)
+2 awa_Deva CentralIndic Awadhi (Devanagari script)
+3 ben_Beng Magadhi Bengali (Bengali script)
+4 bho_Deva Magadhi Bhojpuri (Devanagari script)
+5 bra_Deva CentralIndic Braj Bhasha (Devanagari script)
+6 brx_Deva TibetoBurman Bodo (Devanagari script)
+7 doi_Deva WesternIndic Dogri (Devanagari script)
+8 eng_Latn WestGermanic English (Latin script)
+9 gom_Deva Maharashtri Goan Konkani (Devanagari script)
+10 gon_Deva Dravidian Gondi (Devanagari script)
+11 guj_Gujr WesternIndic Gujarati (Gujarati script)
+12 hin_Deva CentralIndic Hindi (Devanagari script)
+13 hingh_Deva CentralIndic Hinglish (Devanagari script)
+14 hoc_Wara AustroAsiatic Ho (Warang Citi script)
+15 kan_Knda Dravidian Kannada (Kannada script)
+16 kas_Arab WesternIndic Kashmiri (Arabic script)
+17 kas_Deva WesternIndic Kashmiri (Devanagari script)
+18 kha_Latn AustroAsiatic Khasi (Latin script)
+19 lus_Latn TibetoBurman Mizo (Latin script)
+20 mag_Deva Magadhi Magahi (Devanagari script)
+21 mai_Deva Magadhi Maithili (Devanagari script)
+22 mal_Mlym Dravidian Malayalam (Malayalam script)
+23 mar_Deva Maharashtri Marathi (Devanagari script)
+24 mni_Beng TibetoBurman Manipuri (Bengali script)
+25 mni_Mtei TibetoBurman Meitei (Meitei script)
+26 npi_Deva CentralIndic Nepali (Devanagari script)
+27 ory_Orya Magadhi Odia (Odia script)
+28 pan_Guru WesternIndic Punjabi (Gurmukhi script)
+29 san_Deva Vedic Sanskrit (Devanagari script)
+30 sat_Olck AustroAsiatic Santali (Ol Chiki script)
+31 sin_Sinh Maharashtri Sinhala (Sinhala script)
+32 snd_Arab WesternIndic Sindhi (Arabic script)
+33 snd_Deva WesternIndic Sindhi (Devanagari script)
+34 tam_Taml Dravidian Tamil (Tamil script)
+35 tcy_Knda Dravidian Tulu (Kannada script)
+36 tel_Telu Dravidian Telugu (Telugu script)
+37 urd_Arab CentralIndic Urdu (Arabic script)
+38 xnr_Deva CentralIndic Kangri (Devanagari script)
+8.2 Language Family and Group
+We utilized language families and geographical proximity information based on linguistic similarities such as grammar, phonology, and syntax to model the mentioned translation and translation-related tasks.
+Geographical proximity and cultural exchanges create additional linguistic overlaps beyond family and group classifications.
+For example, Hindi and Marathi share vocabulary due to geographic closeness, while Tamil and Telugu exhibit shared phonetic and grammatical patterns due to their Dravidian roots.
+For the languages considered, key families and their groups in the Indian subcontinent include:
+  • Indo-European [^38] https://en.wikipedia.org/wiki/Indo-European˙languages: This family includes various groups such as Central Indic, Western Indic, Magadhi, Maharashtri, and Vedic, encompassing languages with shared syntactic and lexical roots.
+These commonalities enable efficient shared representations in multilingual models.
+    • Magadhi: Includes Assamese (Bengali script), Bengali (Bengali script), Bhojpuri (Devanagari script), Magahi (Devanagari script), Maithili (Devanagari script), and Odia (Odia script), characterized by geographic proximity and linguistic similarities.
+    • Western Indic: Includes Dogri (Devanagari script), Gujarati (Gujarati script), Kashmiri (Arabic and Devanagari scripts), Punjabi (Gurmukhi script), and Sindhi (Arabic and Devanagari scripts), reflecting shared historical and cultural influences.
+    • Central Indic: Includes Awadhi (Devanagari script), Braj Bhasha (Devanagari script), Hindi (Devanagari script), Hinglish (Devanagari script), Kangri (Devanagari script), Nepali (Devanagari script), and Urdu (Arabic script), sharing grammatical and lexical roots.
+    • Maharashtri: Includes Marathi (Devanagari script), Sinhala (Sinhala script), and Goan Konkani (Devanagari script), highlighting regional linguistic ties.
+    • Vedic: Includes Sanskrit (Devanagari script), a classical language foundational to many modern Indo-European languages in India.
+  • Dravidian [^39] https://en.wikipedia.org/wiki/Dravidian˙languages: Includes Kannada (Kannada script), Malayalam (Malayalam script), Tamil (Tamil script), Telugu (Telugu script), Tulu (Kannada script), and Gondi (Devanagari script).
+These languages are marked by agglutinative morphology and shared structural features.
+  • Tibeto-Burman [^40] https://en.wikipedia.org/wiki/Tibeto-Burman˙languages: Includes Bodo (Devanagari script), Manipuri (Bengali script), Meitei (Meitei script), and Mizo (Latin script), characterized by structural similarities despite script diversity.
+  • Austro-Asiatic [^41] https://en.wikipedia.org/wiki/Austroasiatic˙languages: Includes Khasi (Latin script), Santali (Ol Chiki script), and Ho (Warang Citi script), each with distinct phonetic and grammatical features.
+  • West Germanic: English (Latin script), often serving as a bridge language in multilingual models.
+In computational systems, these relationships support transfer learning.
+Shared features among Indic languages like Hindi, Bengali, and Marathi improve translation quality for low-resource pairs.
+Similarly, shared morphology in Dravidian languages enhances generalization across Tamil, Telugu, and Kannada.
+To uniquely identify languages as presented in Table 9, we used a tuple structure combining the language family, script, and a three-character code.
+For instance, Hindi is represented as CentralIndic+hin_Deva, where CentralIndic denotes the family, hin is the language code, and Deva identifies the Devanagari script.
+8.3 Indian Language specific Tokenizer
+We implemented a SentencePieceProcessor-based [^42] https://github.com/google/sentencepiece subword model for 35 Indian subcontinent languages and English, using a vocabulary size of 48,000 tokens.
+To train the tokenizer, we randomly mixed 10 million parallel corpora from each language, ensuring robust coverage of script-specific subwords.
+This approach captures the rich morphological and orthographic variations of Indian languages while maintaining compatibility with English.
+The shared vocabulary, covering 99.99% of characters across languages, facilitates multilingual learning, improves translation quality for low-resource pairs, and leverages linguistic similarities among Indian languages, all while ensuring computational efficiency.
+9 Domains
+For machine translation and related applications, we focused on education, law, and healthcare as major domains.
+The education domain spans diverse disciplines, including Natural Science, Mathematics, Law, and Computer Science, reflecting demand in scientific, legal, and technological communication.
+Additional fields like Teaching Methods, Psychology, Political Science, Management, Tourism, History, Communication Skills, Economics, Marketing, Textiles, and Health highlight the interdisciplinary nature of educational content.
+General domains further demonstrate the broad applicability of translation tools for educational and administrative purposes, providing a strong foundation for multilingual education and knowledge dissemination.
+10 Modeling
+BhashaVerse goes beyond traditional translation models by incorporating the aforementioned tasks such as error identification, automatic post-editing, and grammar correction, significantly enhancing translation quality.
+It is designed to handle the complexities of Indian languages with a medium-range parameter encoder-decoder (Vaswani et al., 2017) model.
+The model is versatile, supporting fine-tuning and adaptability to newer domains and tasks.
+Optimized for smaller GPUs, BhashaVerse delivers state-of-the-art performance while remaining accessible for systems with limited computational resources.
+10.1 Input and Output structure for Multi Task Learning
+In our learning setup, we utilize multi-task multi-lingual corpora to minimize reliance on additional data.
+For sequence-to-sequence training, the objective is configured to address current and future tasks.
+Inspired by structure-driven natural language tasks, we use a JSON-based input-output format (Béchard and Ayala, 2024), designed to support current applications and adapt to future use cases.
+Below is the defined JSON structure:
+Input JSON:
+Ψ"task": "task and source to target language information",
+Ψ"domain": "",
+Ψ"input": { input details based on the task }
+Output JSON:
+Ψ"task": "task and source to target language information",
+Ψ"domain": "",
+Ψ"output": { output details based on the task }
+  • task: Represents the type of task being performed, such as translation or other language-related tasks.
+It also specifies the source and target language information.
+For example, it could be a translation task from English to Hindi or a translation post-editing task from English to Hindi, etc.
+This provides the model with clear task context.
+  • domain: Indicates the domain of the task, such as general, medical, or legal.
+This key allows the model to adapt its behavior based on domain-specific requirements.
+In cases where domain information is not available, then the value would be general.
+  • input: A nested JSON object containing the language information and its corresponding input text for processing.
+  • output: A nested JSON object containing the output language information and the processed output text.
+This JSON structure ensures clarity and modularity in handling inputs and outputs across tasks.
+By explicitly defining the task and domain in both the input and output, the model can effectively manage multiple tasks, while the standardized format supports scalability and integration into various applications for new tasks.
+For this work, we define the task, input, and output sample examples for each of our considered tasks as follows.
+10.1.1 Machine Translation
+Input Json
+"task": "Translation
+        \$WestGermanic+eng_Latn#CentralIndic+hin_Deva",
+"domain": "general",
+"input": {"WestGermanic+eng_Latn": "This committee, the whole
+            state, again, where should be the capital?"},
+Output Json
+"task": "Translation
+         \$WestGermanic+eng_Latn#CentralIndic+hin_Deva",
+"domain": "general",
+"output": {"CentralIndic+hin_Latn": "ye samiti, poora pradesh,
+            phir kahaan ho raajadhaanee?"}
+For the task of translation with English-to-Hindi as an example, the input data format is a JSON object specifying the task as translation between the source language (English, WestGermanic+eng_Latn) and the target language (Hindi, CentralIndic+hin_Latn).
+The domain of the text is set as "general," and the input field contains the English sentence, such as "This committee, the whole state, again, where should be the capital?".
+The corresponding output JSON contains the Hindi translation in the output field, e.g., "ye samiti, poora pradesh, phir kahaan ho raajadhaanee?".
+This format trains the model to handle multilingual translation tasks, where the encoder processes the source language text and the decoder generates the target language text.
+10.1.2 Grammar Correction
+Input Json
+"task": "Correction\$Incorrect
+         WestGermanic+eng_Latn#WestGermanic+eng_Latn",
+"domain": "general",
+"input": {"Incorrect
+           WestGermanic+eng_Latn": "This committee, the whole
+           state,
+           again, whrn shuld be the capital?"},
+Output Json
+"task": "Correction\$Incorrect
+         WestGermanic+eng_Latn#WestGermanic+eng_Latn",
+"domain": "general",
+"output": {"Corrected
+            WestGermanic+eng_Latn": "This committee, the whole
+            state,
+            again, where should be the capital?"},
+In the grammar correction task, the input JSON specifies the task as correcting grammatically incorrect text for a language.
+For instance, the input contains an English sentence with errors, such as "This committee, the whole state, again, whrn shuld be the capital?", tagged as Incorrect WestGermanic+eng_Latn.
+The output JSON includes the corrected text in the output field, e.g., "This committee, the whole state, again, where should be the capital?".
+This task trains the model to map ungrammatical inputs to corrected outputs, helping to refine text for improved quality and clarity.
+10.1.3 Machine Translation Post-editing
+Input Json
+"task": "Translation post editing
+Ψ \$WestGermanic+eng_Latn#Dravidian+tel_Latn",
+"domain": "Computer science",
+"input": {"WestGermanic+eng_Latn": "Light contrast contributes
+          to this kind of
+          animation and makes it more impressive.",
+          "Dravidian+tel_Latn": "Tlikapi krs
+          rakamaina
+          ynimanku dhadam
+          cstundi mariyu idi marinta kaukuundi."}
+Output Json
+"task": "Translation post editing
+Ψ \$WestGermanic+eng_Latn#Dravidian+tel_Latn",
+"domain": "Computer science",
+"output": {"post edited Dravidian+tel_Latn": "Lai krs
+            rakamaina ynimanku dhadam
+            cstundi mariyu dnini marinta
+            kaukunl cstundi."}
+This task involves refining machine-translated text for improved accuracy and fluency.
+For example, for a general domain post-editing task, the input JSON includes the source text in English, the machine-generated translation in Telugu, and the task type as Translation post editing.
+For example, the English input might be "Light contrast contributes to this kind of animation and makes it more impressive." with a machine-generated Telugu translation "Tēlikapāṭi kāṇṭrāsṭ ī rakamaina yānimēṣanku dōhadam cēstundi mariyu idi marinta ākaṭṭuṇṭundi." The output JSON provides the post-edited Telugu translation, e.g., "Laiṭ kāṇṭrāsṭ ī rakamaina yānimēṣanku dōhadam cēstundi mariyu dānini marinta ākaṭṭukunēlā cēstundi." This data trains the model to refine translations, ensuring better alignment with linguistic norms and domain-specific requirements.
+10.1.4 Machine Translation Direct Assessment
+Input Json
+"task": "Translation direct assessment
+Ψ \$WestGermanic+eng_Latn#Dravidian+tel_Latn",
+"domain": "Computer science",
+"input": {"WestGermanic+eng_Latn": "Light contrast contributes
+          to this kind of animation and makes
+          it more impressive.",
+Ψ   "gold Dravidian+tel_Latn": "Lai krs  rakamaina
+          ynimanku dhadam cstundi mariyu dnini
+          marinta kaukunl cstundi."
+Ψ   "system Dravidian+tel_Latn": "Tlikapi krs
+          rakamaina ynimanku dhadam cstundi mariyu
+          idi marinta kaukuundi."}
+Output Json
+"task": "Translation direct assessment
+Ψ \$WestGermanic+eng_Latn#Dravidian+tel_Latn",
+"domain": "Computer science",
+"output": {"direct assessment score out of 100": 78.34184}
+In this task, the model is trained to evaluate the quality of a translation against a gold-standard reference.
+For example, the input JSON includes the source English text, the gold-standard Telugu translation, and the system-generated Telugu translation.
+For example, the input text might be "Light contrast contributes to this kind of animation and makes it more impressive." with a gold translation and a system-generated translation for comparison.
+The output JSON provides a direct assessment score, such as 78.34, reflecting the quality of the system output relative to the reference.
+This trains the model to assign quality scores based on reference-based evaluation metrics.
+10.1.5 Machine Translation Quality Estimation
+Input Json
+"task": "Translation quality estimation
+Ψ \$WestGermanic+eng_Latn#Dravidian+tel_Latn",
+"domain": "Computer science",
+"input": {"WestGermanic+eng_Latn": "Light contrast contributes
+            to this kind of animation and makes
+            it more impressive.",
+        "Dravidian+tel_Latn": "Tlikapi krs  rakamaina
+        ynimanku dhadam cstundi mariyu
+        idi marinta kaukuundi."}
+Output Json
+"task": "Translation quality estimation
+Ψ \$WestGermanic+eng_Latn#Dravidian+tel_Latn",
+"domain": "Computer science",
+"output": {"quality estimation score out of 100": 78.34184}
+This task evaluates translation quality without a gold-standard reference.
+For example, the input JSON includes the English source text and its Telugu translation, with the task defined as Translation quality estimation and domain as Computer Science.
+For example, the input might contain the English sentence "Light contrast contributes to this kind of animation and makes it more impressive." and the Telugu translation is "Tēlikapāṭi kāṇṭrāsṭ ī rakamaina yānimēṣanku dōhadam cēstundi mariyu idi marinta ākaṭṭukuṇṭundi." The output JSON provides a quality score, e.g., 78.34.
+This format supports training for quality estimation tasks in scenarios where reference translations are unavailable.
+10.1.6 Machine Translation Error Identification
+Input Json
+"task": "Translation error marking
+Ψ \$WestGermanic+eng_Latn#Dravidian+tel_Latn",
+"domain": "Computer science",
+"input": {"WestGermanic+eng_Latn": "Light contrast contributes
+           to this kind of animation
+           and makes it more impressive.",
+Ψ   "Dravidian+tel_Latn": "Tlikapi krs  rakamaina
+           ynimanku dhadam cstundi mariyu
+            idi marinta kaukuundi."}
+Output Json
+"task": "Translation error marking
+Ψ \$WestGermanic+eng_Latn#Dravidian+tel_Latn",
+"domain": "Computer science",
+"output": {"error marked Dravidian+tel_Latn":
+            "<e>tlikapi</e> krs
+             rakamaina ynimanku <e>dhadam</e> cstundi
+            <e>mariyu</e> <e>idi</e> marinta
+            <e>kaukuundi.</e>"}
+The error identification task identifies and highlights errors in translated text.
+Again as an example, the input JSON specifies the source English text and its Telugu translation, with the task labeled as Translation error marking for the domain computer science.
+For example, the input might include "Light contrast contributes to this kind of animation and makes it more impressive." and the Telugu translation "Tēlikapāṭi kāṇṭrāsṭ ī rakamaina yānimēṣanku dōhadam cēstundi mariyu idi marinta ākaṭṭuṇṭundi." The output JSON highlights errors using tags like <e>.
+For instance, "<E>tēlikapāṭi</e> kāṇṭrāsṭ <e>dōhadam</e> cēstundi <e>marinta</e> ākaṭṭuṇṭundi." This trains the model to detect and mark specific areas needing improvement in translations.
+10.1.7 Machine Translation Error Identification and Correction
+Input Json
+"task": "Translation error marking and correction
+Ψ \$WestGermanic+eng_Latn#Dravidian+tel_Latn",
+"domain": "Computer science",
+"input": {"WestGermanic+eng_Latn": "Light contrast contributes
+          to this kind of animation and
+          makes it more impressive.",
+Ψ   "Dravidian+tel_Latn": "Tlikapi krs
+          rakamaina ynimanku dhadam cstundi
+          mariyu idi marinta kaukuundi."}
+Output Json
+"task": "Translation error marking and correction
+Ψ \$WestGermanic+eng_Latn#Dravidian+tel_Latn",
+"domain": "Computer science",
+"output": {"error marked Dravidian+tel_Latn": "<e>tlikapi</e>
+            krs  rakamaina ynimanku <e>dhadam</e>
+            cstundi <e>mariyu</e> <e>idi</e> marinta
+            <e>kaukuundi.</e>",
+Ψ   "post editedDravidian+tel_Latn": "Lai krs
+           rakamaina ynimanku dhadam cstundi mariyu
+           dnini marinta kaukunl cstundi."}
+This task combines error marking with automatic correction.
+The input JSON includes the source text, its corresponding translation, and the defined task as Translation error marking and correction.
+For example, the English text "Light contrast contributes to this kind of animation and makes it more impressive." and the Telugu translation "Tēlikapāṭi kāṇṭrāsṭ ī rakamaina yānimēṣanku dōhadam cēstundi mariyu idi marinta ākaṭṭuṇṭundi." are provided.
+The output JSON contains both the error-marked translation (e.g., "<e>tēlikapāṭi</e> kāṇṭrāsṭ <e>dōhadam</e> cēstundi <e>marinta</e> ākaṭṭuṇṭundi.") and the corrected text (e.g., "Laiṭ kāṇṭrāsṭ ī rakamaina yānimēṣanku dōhadam cēstundi mariyu dānini marinta ākaṭṭukunēlā cēstundi.").
+This format trains the model to not only identify errors but also generate refined, corrected translations.
+10.2 Architecture
+We use an encoder-decoder model using the fairseq toolkit [^43] https://github.com/facebookresearch/fairseq for its ability to handle complex language structures in bilingual and multilingual contexts.
+The encoder encodes input sentences with positional embeddings, capturing key features and context using self-attention.
+The decoder generates translations word by word, leveraging self-attention to focus on relevant input parts and maintain coherence with previously generated words (Vaswani et al., 2017).
+The transformer encoder-decoder models are trained from scratch using the parameter configurations below.
+  • Input: subword tokens.
+  • Embedding Size: 1024.
+  • Feedforward Size: 8192.
+  • Layers: Encoder: 18, Decoder: 18.
+  • Attention Heads: 16.
+  • Dropout: 0.30.
+  • Max Word Sequence Length: 1200.
+  • Batch Size: 1200 tokens.
+  • Initial Learning Rate: 2e-5.
+  • Optimizer: Adam.
+  • Label Smoothing: 0.1.
+  • Precision: 16-bit floating point.
+  • Early Stopping: Triggered if no increase in training loss for 10 epochs.
+  • Beam Size: 15.
+11 Training Process
+The model was trained in two stages using 32 A100 GPUs, each with 40 GB of memory.
+In the first stage, trained on all available corpora (mentioned in above sections) across specified domains and languages for the considered tasks with JSON format.
+In the second stage, as a continual training, we focused on exclusive training using human-developed corpora from various sources of different tasks, so that the model would produce outputs that are more aligned with human-like language generation.
+12 Evaluation
+In this section, we present the evaluation of machine translation and related tasks performed on our trained model.
+These tasks include machine translation, grammar correction, post-editing, human direct assessment of translations, quality assessment, error identification, and combined error identification and correction.
+For machine translation, we relied on existing benchmark corpora such as Flores (Goyal et al., 2021) and IN22 (Gala et al., 2023), which provide robust datasets for assessing translation performance across diverse language pairs.
+However, for the other tasks, we could not find suitable benchmark datasets that covered all the language pairs included in this work.
+Consequently, we utilized our reserved development corpora (1,000 instances per task per language/pair), specifically removed and curated from the training data, for these tasks.
+Table 10 outlines the sources and sizes of the corpora used for each task, along with the corresponding evaluation metrics applied to assess model performance.
+This approach ensures a comprehensive evaluation, particularly for less-studied language pairs and specialized tasks.
+We discuss the average result for each task as follows.
+Task Benchmark Source Language Pairs
+Size
+Metric Average Score
+Machine Translation FLORES + IN22 36 * 36 Languages (English and 35 Indian subcontinent Languages) 517638 + 384560 BLEU 25.45
+CHAF3 53.81
+COMET-22 0.8138
+Grammar Correction Development Corpora
+22 Schedule
+Indian Languages
+22000 BLEU 90.55
+CHRF 97.61
+Machine Translation
+Post-editing
+Development Corpora
+English to Hindi, Kannada,
+Marathi, Telugu, Bangla
+5000 BLEU 84.26
+CHAF3 95.47
+Machine Translation
+Direct Assessment
+Development Corpora
+English to Hindi, Bengali,
+Tamil, Telugu, Gujarati,
+Marathi, Kannada, Malayalam,
+Assamese, Punjabi, and Urdu
+11000
+Spearman
+Co-relation
+0.46202
+Machine Translation
+Quality Estimation
+Development Corpora
+English to Hindi, Bengali,
+Tamil, Telugu, Gujarati,
+Marathi, Kannada, Malayalam,
+Assamese, Punjabi, and Urdu
+11000
+Spearman
+Co-relation
+0.47482
+Machine Translation
+Error Identification
+Development Corpora
+English to Hindi, Kannada,
+Marathi, Telugu, Bangla
+5000 F1 Score 41.23
+Machine Translation Error Identification and Correction Development Corpora English to Hindi, Kannada, Marathi, Telugu, Bangla
+Error Identification
+5000
+F1 Score 39.23
+Correction 5000 BLEU 84.02
+CHAF3 94.77
+Table 10: Performance evaluation of the multi-task, multilingual sequence-to-sequence model with 2 billion parameters across diverse tasks.
+The table presents the benchmark sources, language pairs, dataset sizes, evaluation metrics, and average scores for machine translation, grammar correction, post-editing, quality assessment, and error identification and correction.
+The model is evaluated on machine translation using benchmark corpora such as FLORES and IN22, covering available languages, including English and other Indian subcontinent languages.
+With a combined dataset size of 902,198 sentences, the model achieved an average BLEU score (Papineni et al., 2002) of 25.45, CHRF3 score (Popović, 2015) of 53.81, and COMET-22 score (Rei et al., 2022a) of 0.8138.
+These results reflect the model’s capacity to deliver high-quality translations while effectively handling linguistic diversity and complexity.
+Multi-task learning significantly enhanced the model’s performance across related tasks.
+For grammar correction, on 22,000 grammatically incorrect sentences, the model achieved an average BLEU and CHRF3 scores of 90.55 and 97.61, respectively.
+This highlights the model’s ability to refine grammatical accuracy while benefiting from its translation training.
+Similarly, post-editing tasks involving English-to-Hindi, Kannada, Marathi, Telugu, and Bangla translations demonstrated high scores, achieving an average BLEU score of 84.26 and a CHRF3 score of 95.47, showing how translation training supports fine-tuning and error reduction in post-editing.
+Tasks such as direct assessment and quality estimation, evaluated on 11 Indian languages using 11,000 sentences, further underscore the effectiveness of multi-task learning.
+The Spearman correlation scores of 0.46202 for direct assessment and 0.47482 for quality estimation reveal the model’s alignment with human judgments and its ability to predict translation quality, aided by shared learning across tasks.
+Error identification and correction tasks, evaluated on five language pairs (English-to-Hindi, Kannada, Marathi, Telugu, and Bangla), demonstrated the model’s robustness in identifying noisy translation chunks.
+Error identification achieved an F1 score of 41.23, while the combined error identification and correction task yielded an average BLEU (84.02) and CHRF3 (94.77) scores for corrected outputs.
+From the numbers, we can argue that the model effectively handles tasks such as machine translation, grammar correction, post-editing, quality assessment, and error identification and correction across multiple Indian languages.
+The results demonstrate its superior translation performance and ability to generalize across tasks, highlighting the benefits of multi-task learning.
+13 Conclusion
+This work explores the capabilities of a 2-billion-parameter multilingual, multi-task sequence-to-sequence model in addressing the challenges of translating Indian languages.
+Through multi-task learning, the model excels in tasks such as machine translation, grammar correction, post-editing, quality estimation, and error identification, effectively handling the syntactic, morphological, and script complexities of Indian languages.
+The study emphasizes the value of curated datasets, domain-specific corpora, and robust evaluation frameworks in overcoming low-resource language challenges.
+The model demonstrates scalability across 36 × 36 language pairs, supports domain-specific translations, and ensures contextual coherence in discourse-level tasks, showcasing its versatility.
+This research, created 10 billion parallel corpora along with millions of related task corpora across 36 Indian subcontinent languages and a trained model, shows the potential to bridge linguistic gaps and support multilingual communication in critical areas such as healthcare, education, and governance.
+Future work will focus on expanding to additional low-resource languages, enhancing discourse-level performance, and refining automatic post-editing systems, paving the way for advancements in machine translation within resource-constrained settings.